### 05 Mar 2024 v1

# Locating Superspreaders in SI Cascades

Locating Superspreaders in SI Cascades
Elchanan Mossel MIT; elmos@mit.edu    Anirudh Sridhar MIT; anisri@mit.edu
(March 5, 2024)
1 Overview
  • Define a class of graphs where almost all vertices have degree at most $d$, but a few (constant number) might have degree at least $D$.
  • Cascade starts at either a designated vertex (which we have no control over), or a uniform random vertex
  • Observe the sequence of infection times for $K$ samples. What is the min value of $K$ that will let us accurtely determine which nodes have degree at least $D$?
  • If $D\sim n^{\alpha}$ for $\alpha\in(2/3,1)$, then $K=3$ suffices! And from a single sample, we can identify when a super-spreading event happens.
  • If $\alpha\in(0,1/2)$, $K=\log(n)/\log\log(n)$ samples is necessary and sufficient. This is comparable to the number of samples needed to learn the entire network!
  • Open regime: What is the right sample complexity for $\alpha\in(1/2,2/3)$?
2 Todos
  • Write introduction, and format according to COLT requirements
  • Finish proof of achievability when $G$ is a tree and $\alpha\in(0,1)$
  • Figures: Visualizing graphs for lower bounds as well as intuition behind the second derivative method
  • Simulations(?): I think a good synthetic graph would be a binary tree with a high-degree leaf, but this may feel too synthetic for readers. It could be a good illustrative example, though, where a high-degree node early on is readily detectable from the infection curve, but a high-degree leaf requires the second derivative analysis to detect it.
    Another possibility is to find a decent enough network with a heavy-tailed degree distribution, and generate samples based on that causal network.
3 Notation
  • For a graph $G$ and a starting location $v_{0}$, let $\mathcal{SI}(G,v_{0})$ be the set of infection times generated by an SI process on $G$ starting at $v_{0}$.
4 Cascade dynamics
Let $G$ be an undirected graph and let $\lambda\geq 0$ denote the pairwise interaction rate between any two vertices in $G$ connected by an edge. For a time index $t\geq 0$, we denote $I[0,t]$ to be the set of vertices infected in the interval $[0,t]$. Given $I[0,t]$, the cascade evolves as follows: for $v\in V\setminus I[0,t]$, it holds that as $\epsilon\to 0$,
$\operatorname{\mathbb{P}}\left(\left.v\in I[0,t+\epsilon]\right|I[0,t]\right)=\epsilon\lambda|{\mathcal{N}}(v)\cap I[0,t]|+o(\epsilon).$ (4.1)
For a given vertex $v\in V$, we formally define its infection time in the SI model to be
$T(v):=\inf\{t\geq 0:v\in I[0,t]\}.$
It is well-known that the dynamics (4.1) is equivalent to First Passage Percolation on $G$ with $\mathrm{Exp}(\lambda)$ edge weights, which leads to an extremely useful representation of the SI process.
In a bit more detail, let $\{{\mathcal{F}}_{s}\}_{s\geq 0}$ denote the natural filtration corresponding to the SI process. Fix $t\geq 0$, and let $\{F(e)\}_{e\in E}$ be a collection of i.i.d. $\mathrm{Exp}(\lambda)$ random variables that are independent of ${\mathcal{F}}_{t}$. For a path [^1] Recall that a path is a finite sequence of edges $e_{1},\ldots,e_{m}$ such that for each $i\geq 1$, $e_{i}$ and $e_{i+1}$ share exactly one endpoint. $P$ in the graph, we define the weight of the path to be $\mathsf{weight}(P):=\sum_{e\in P}F(e)$. Then we have the following result.
Proposition 4.1.
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Suppose that $v\in V\setminus{\mathcal{I}}[0,t]$. Let ${\mathcal{P}}(v,t)$ be the collection of paths from $I[0,t]$ to $v$ where only the first edge in the path has an endpoint in $I[0,t]$. Then the conditional distribution of $T(v)$ with respect to ${\mathcal{F}}_{t}$ is given by
$T(v)=t+\min_{P\in{\mathcal{P}}(v,t)}\mathsf{weight}(P).$
This result is standard, and essentially follows from the memoryless property of exponential distributions. For completeness, we provide the proof in Appendix A. [say some further remarks about FPP, this is a bit abrupt]
Using the representation in Proposition 4.1, we can derive a number of useful results on the infection times of vertices.
Lemma 4.2.
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Suppose that $v\in V\setminus{\mathcal{I}}[0,t]$. Then $T(v)-t$ is stochastically dominated by a $\mathrm{Exp}(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|)$ random variable.
Proof.
Conditioned on ${\mathcal{F}}_{t}$, Proposition 4.1 implies that
$T(v)-t$ $=\min_{P\in{\mathcal{P}}(v,t)}\mathsf{weight}(P)$
$\leq\min_{P\in{\mathcal{P}}(v,t):|P|=1}\mathsf{weight}(P)$
$=\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v).$
Due to properties of exponential random variables, we have that
$\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v)\sim\mathrm{Exp}\left(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\right),$
which proves the claimed result. ∎
[lower bound on infection time for trees]
Lemma 4.3.
Suppose that $G$ is a tree. Let $t\geq 0$, condition on ${\mathcal{F}}_{t}$, and suppose that $v\in V\setminus{\mathcal{I}}[0,t]$. Then $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda)$ random variable.
Proof.
Since $G$ is a tree, there is a unique path from the infection source $s$ to $v$. As a result, $v$ can never have more than one infected neighbor, so $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda)$ random variable. ∎
[lower bound on infection time]
Lemma 4.4.
Let $t\geq 0$, condition on ${\mathcal{F}}_{t}$, and suppose that $v\in V\setminus{\mathcal{I}}[0,t]$. Then $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda|{\mathcal{N}}(v)|)$ random variable.
Proof.
Conditioned on ${\mathcal{F}}_{t}$, we have that
$T(v)-t=\min_{P\in{\mathcal{P}}(v,t)}\mathsf{weight}(P)\geq\min_{u\in{\mathcal{N}}(v)}F(u,v),$
where the inequality follows since any path ending in $v$ must include one of the edges incident to $v$. The claim follows since $\min_{u\in{\mathcal{N}}(v)}F(u,v)\sim\mathrm{Exp}(\lambda|{\mathcal{N}}(v)|)$. ∎
[number of infected neighbors of high-deg vertex by the time it gets infected]
We next define a useful event concerning the number of infected neighbors of high-degree vertices.
[potentially change to $n^{1/2+\epsilon}$, or similar. It would look cosmetically simpler and probs would be easier to digest.]
Definition 4.5.
The event ${\mathcal{A}}$ holds if and only if
$|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,T(v)]|\leq 2d\sqrt{\deg(v)}\log^{2}n,$
for all $v\in V$ such that $\deg(v)\geq D$.
Lemma 4.6.
It holds that $\operatorname{\mathbb{P}}({\mathcal{A}})=1-o(1)$.
Proof.
Let $\deg(v)\geq D$, and define the event
${\mathcal{E}}_{v}:=\left\{|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,T(v)]|\geq 2d\sqrt{\deg(v)}\log^{2}n\right\}.$
Let $T$ be the (random) time at which $\sqrt{\deg(v)}\log n$ neighbors of $v$ are first infected.
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Let us also assume that $T=t$ (an event that is ${\mathcal{F}}_{t}$-measurable), so that $|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|=\sqrt{\deg(v)}\log n$. If $T(v)\leq t$, then we have that $\operatorname{\mathbb{P}}({\mathcal{A}}|{\mathcal{F}}_{t})=0$. Otherwise, Lemma 4.2 implies that $T(v)-t$ is stochastically dominated by a $\mathrm{Exp}(\lambda\sqrt{\deg(v)}\log n)$ random variable. Hence
$\operatorname{\mathbb{P}}\left(\left.T(v)\geq t+\frac{1}{\lambda\sqrt{\deg(v)}}\right|{\mathcal{F}}_{t}\right)\leq\frac{1}{n}.$ (4.2)
Next, we upper bound the number of neighbors of $v$ that become infected in the interval $[t,t+1/(\lambda\sqrt{\deg(v)})]$. By our assumption on the distance between high-degree vertices [ref somewhere], all neighbors of $v$ have degree at most $d$. Lemma 4.4 therefore implies that for $u\in{\mathcal{N}}(v)\setminus{\mathcal{I}}[0,t]$, $T(u)-t$ stochastically dominates a $\mathrm{Exp}(\lambda d)$ random variable, and
$\operatorname{\mathbb{P}}\left(\left.T(u)-t\leq\frac{1}{\lambda\sqrt{\deg(v)}}\right|{\mathcal{F}}_{t}\right)\leq 1-e^{-d/\sqrt{\deg(v)}}\leq\frac{d}{\sqrt{\deg(v)}}.$
It follows that
$\operatorname{\mathbb{E}}\left[\left.|{\mathcal{N}}(v)\cap{\mathcal{I}}[t,T(v)]|\mathbf{1}\left(T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right)\right|{\mathcal{F}}_{t}\right]$ $\leq\operatorname{\mathbb{E}}\left[\left.\left|{\mathcal{N}}(v)\cap{\mathcal{I}}\left[t,t+\frac{1}{\lambda\sqrt{\deg(v)}}\right]\right|\right|{\mathcal{F}}_{t}\right]$
$=\sum_{u\in{\mathcal{N}}(v)\setminus{\mathcal{I}}[0,t]}\operatorname{\mathbb{P}}\left(\left.T(u)\leq t+\frac{1}{\lambda\sqrt{\deg(v)}}\right|{\mathcal{F}}_{t}\right)$
$\leq d\sqrt{\deg(v)}.$
As a consequence,
$\operatorname{\mathbb{E}}\left[\left.\left|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,T(v)]\right|\mathbf{1}\left(T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right)\right|{\mathcal{F}}_{t}\right]$
$\hskip 28.45274pt\leq|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|+\operatorname{\mathbb{E}}\left[\left.|{\mathcal{N}}(v)\cap{\mathcal{I}}[t,T(v)]|\mathbf{1}\left(T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right)\right|{\mathcal{F}}_{t}\right]$
$\hskip 28.45274pt\leq\sqrt{\deg(v)}\log n+d\sqrt{\deg(v)}\leq 2d\sqrt{\deg(v)}\log n,$
where, in the final expression, we have used our assumption $T=t$. Next, Markov’s inequality implies
$\operatorname{\mathbb{P}}\left(\left.{\mathcal{E}}_{v}\cap\left\{T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right\}\right|{\mathcal{F}}_{t}\right)\leq\frac{1}{\log^{2}n}.$
A union bound along with (4.2) shows that
$\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{t})\leq\operatorname{\mathbb{P}}\left(\left.{\mathcal{E}}_{v}\cap\left\{T(v)<t+\frac{1}{\lambda\sqrt{\deg(v)}}\right\}\right|{\mathcal{F}}_{t}\right)+\operatorname{\mathbb{P}}\left(\left.T(v)\geq t+\frac{1}{\lambda\sqrt{\deg(v)}}\right|{\mathcal{F}}_{t}\right)\leq\frac{2}{\log^{2}n}.$ (4.3)
Next, since $T<\infty$ almost surely (since all vertices in the graph become infected in finite time), it holds that
[…]
which, in light of (4.3), implies that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{T})\leq 2/\log^{2}(n)$. It follows that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v})=\operatorname{\mathbb{E}}[\operatorname{\mathbb{P}}({\mathcal{E}}_{v}|{\mathcal{F}}_{T})]\leq 2/\log^{2}(n)$. Finally, we note that ${\mathcal{A}}^{c}:=\bigcup_{v\in G:\deg(v)\geq D}{\mathcal{E}}_{v}$, hence a union bound shows that
$\operatorname{\mathbb{P}}({\mathcal{A}}^{c})\leq\sum_{v\in G:\deg(v)\geq D}\operatorname{\mathbb{P}}({\mathcal{E}}_{v})=o(1).$
Lemma 4.7.
Let $v\in V$ and suppose that $\deg(v)\lambda^{2}\delta^{2}\leq 1$. Fix $t\geq 0$, and let $v\in V\setminus{\mathcal{I}}[0,t]$. Then
($\operatorname{\mathbb{P}}\left(\left.\{T(v)\in[t,t+\delta]\}\cap{\mathcal{E}}\right|{\mathcal{F}}_{t}\right)\leq 3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}.$
[update ${\mathcal{E}}_{t}$ to ${\mathcal{A}}_{t}$]
Proof.
Define the event
${\mathcal{E}}_{t}:=\left\{\text{For all $v\in V$ such that $\deg(v)\geq D$ and$T(v)>t$, it holds that $|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq 2d\sqrt{\deg(v)}\log^{2}n$}\right\}.$
Notice that ${\mathcal{E}}_{t}$ is ${\mathcal{F}}_{t}$-measurable, and ${\mathcal{A}}\subset{\mathcal{E}}_{t}$. In particular,
($\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}|{\mathcal{F}}_{t}\right)\leq\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}_{t}|{\mathcal{F}}_{t}\right)=\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})\mathbf{1}({\mathcal{E}}_{t}).$ (4.4)
We will therefore proceed by bounding $\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})$ under the assumption that ${\mathcal{E}}_{t}$ holds.
Suppose that $\deg(v)\leq d$. Then by Lemma 4.4, $T(v)-t$ is stochastically lower bounded by a $\mathrm{Exp}(\lambda d)$ random variable, so that
$\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})\leq 1-e^{-\delta\lambda d}\leq\delta\lambda d.$ (4.5)
Next, suppose that $\deg(v)\geq D$, and let ${\mathcal{P}}_{2}(v)$ be the set of paths of length 2 starting from $v$. Then
$T(v)-t$ $=\min_{P\in{\mathcal{P}}(v,t)}\mathsf{weight}(P)$ (4.6)
$\geq\min\left\{\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v),\min_{P\in{\mathcal{P}}_{2}(v)}\mathsf{weight}(P)\right\},$ (4.7)
which can be explained as follows. If the minimizing path $P\in{\mathcal{P}}(v,t)$ in (4.6) has length 1, then the minimum is achieved by $F(u,v)$ for some $u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]$. Otherwise, the minimizing path in (4.6) has length at least 2, which must include a path of length 2 starting at $v$. In light of (4.7), we see that
$\operatorname{\mathbb{P}}(T(v)\leq t+\delta|{\mathcal{F}}_{t})$ $\leq\operatorname{\mathbb{P}}\left(\left.\min\left\{\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v),\min_{P\in{\mathcal{P}}_{2}(v)}\mathsf{weight}(P)\right\}\leq\delta\right|{\mathcal{F}}_{t}\right)$
$\leq\sum_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}\operatorname{\mathbb{P}}(F(u,v)\leq\delta|{\mathcal{F}}_{t})+\sum_{P\in{\mathcal{P}}_{2}(v)}\operatorname{\mathbb{P}}(\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\lambda\delta+|{\mathcal{P}}_{2}(v)|\lambda^{2}\delta^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}+d\left(\deg(v)\lambda^{2}\delta^{2}\right)$
$\stackrel{{\scriptstyle(c)}}{{\leq}}3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}},$ (4.8)
where the inequality $(a)$ follows since $F(u,v)\sim\mathrm{Exp}(\lambda)$ and $\mathsf{weight}(P)\sim\mathrm{Gamma}(2,\lambda)$ for $P\in{\mathcal{P}}_{2}(v)$. The inequality $(b)$ uses the bound $|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq 2d\sqrt{\deg(v)}\log^{2}n$ which holds on the event ${\mathcal{E}}_{t}$, and also uses that the number of paths of length 2 starting from $v$ is at most $d\cdot\deg(v)$, since we assume that high-degree vertices are not neighbors in $G$. Finally, $(c)$ holds by our assumption that $\deg(v)\lambda^{2}\delta^{2}\leq 1$. ∎
Lemma 4.7 can be readily extended to the case where $t$ is replaced by a stopping time. The following result provides one useful example of this idea.
Lemma 4.8.
Let $u,v\in V$ and suppose that $\max\{\deg(u),\deg(v)\}\lambda^{2}\delta^{2}\leq 1$. Then
($\operatorname{\mathbb{P}}\left(\{|T(v)-T(u)|\leq\delta\}\cap{\mathcal{E}}\right)\leq 3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}.$
Proof.
Define the stopping time $T:=\min\{T(u),T(v)\}$. Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Assuming that the ${\mathcal{F}}_{t}$-measurable event $\{T=t\}$ holds and, without loss of generality that $T=T(u)$, we have by Lemma 4.7 that
($\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}|{\mathcal{F}}_{t}\right)\mathbf{1}(T=t)\leq\left(3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}\right)\mathbf{1}(T=t).$
Since $T<\infty$ almost surely, we have that
($\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}\right)$ $=\operatorname{\mathbb{E}}\left[\operatorname{\mathbb{P}}\left(\{T(v)\leq t+\delta\}\cap{\mathcal{E}}|{\mathcal{F}}_{T}\right)\right]$
[…]
$\leq 3d\log^{2}(n)\sqrt{\deg(v)\lambda^{2}\delta^{2}}.$
5 Second derivative method for $\alpha\in(2/3,1]$
5.1 First derivative is approximately the size of the cut
Conditioned on ${\mathcal{F}}_{t}$, we partition ${\mathcal{I}}[t,t+\delta]={\mathcal{I}}_{1}\cup{\mathcal{I}}_{2}$, where
${\mathcal{I}}_{1}$ :=$:=\left\{v\in\partial{\mathcal{I}}(t):\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v)\leq\delta\right\}$
${\mathcal{I}}_{2}$ :=$:=\left\{v\in{\mathcal{I}}[t,t+\delta]:P\in{\mathcal{P}}(v,t)\text{ and }\mathsf{weight}(P)\leq\delta\Rightarrow|P|\geq 2\right\}={\mathcal{I}}[t,t+\delta]\setminus{\mathcal{I}}_{1}.$
We proceed by studying the sizes of ${\mathcal{I}}_{1}$ and ${\mathcal{I}}_{2}$ separately. Our first result characterizes $|{\mathcal{I}}_{1}|$, essentially through an application of Bernstein’s inequality.
Lemma 5.1.
Let $G$ be a tree on $n$ vertices, let $\beta>0$ and assume that $\gamma>\max\{(1-\beta)/2,1-2\beta\}$. Then for any $t\geq 0$, it holds that
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\geq n^{\gamma}\right|{\mathcal{F}}_{t}\right)=o\left(\frac{1}{n}\right).$
Proof.
Since $G$ is a tree, we have for each $v\in\partial{\mathcal{I}}(t)$ that $v$ has exactly one infected neighbor at time $t$. Hence $T(v)-t\sim\mathrm{Exp}(\lambda)$, so
$\operatorname{\mathbb{P}}(v\in{\mathcal{I}}_{1}|{\mathcal{F}}_{t})=\operatorname{\mathbb{P}}(T(v)-t\in[0,n^{-\beta}]|{\mathcal{F}}_{t})=1-e^{-\lambda/n^{\beta}}.$
By the conditional independence of infection times in $\partial{\mathcal{I}}(t)$ with respect to ${\mathcal{F}}_{t}$, we have that $|{\mathcal{I}}_{1}|\sim\mathrm{Bin}(|\partial{\mathcal{I}}(t)|,1-e^{-\lambda/n^{\beta}})$. We can then bound
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}|-|\partial{\mathcal{I}}(t)|\left(1-e^{-\lambda/n^{\beta}}\right)\right|\geq\frac{1}{2}n^{\gamma}\right|{\mathcal{F}}_{t}\right)$ $\stackrel{{\scriptstyle(a)}}{{\leq}}2\exp\left(-\frac{n^{2\gamma}/8}{|\partial{\mathcal{I}}(t)|\left(1-e^{-\lambda/n^{\beta}}\right)+n^{\gamma}/6}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2\exp\left(-\frac{n^{2\gamma}/8}{\lambda n^{1-\beta}+n^{\gamma}/6}\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}o\left(\frac{1}{n}\right).$ (5.1)
Above, $(a)$ is due to Bernstein’s inequality, $(b)$ uses the upper bounds $|\partial{\mathcal{I}}(t)|\leq n$ and $1-e^{-\lambda/n^{\beta}}\leq\lambda/n^{\beta}$, and $(c)$ follows since $\gamma>(1-\beta)/2$.
To prove the claim, it remains to study $|\partial{\mathcal{I}}(t)|(1-e^{-\lambda/n^{\beta}})$. We have that
$\left||\partial{\mathcal{I}}(t)|\left(1-e^{-\lambda/n^{\beta}}\right)-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\stackrel{{\scriptstyle(d)}}{{=}}\left|1-e^{-\lambda/n^{\beta}}-\frac{\lambda}{n^{\beta}}\right||\partial{\mathcal{I}}(t)|\stackrel{{\scriptstyle(e)}}{{\leq}}\lambda^{2}n^{1-2\beta}\stackrel{{\scriptstyle(f)}}{{\leq}}\frac{1}{2}n^{\gamma},$ (5.2)
where $(d)$ follows since $|\partial{\mathcal{I}}(t)|=|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|$ in the case where $G$ is a tree [ref this fact], $(e)$ is due to a Taylor approximation and since $|\partial{\mathcal{I}}(t)|\leq n$, and $(f)$ holds for $n$ sufficiently large since $\gamma>1-2\beta$. Together, (5.1) and (5.2) prove the lemma. ∎
The same ideas can be used to characterize $|{\mathcal{I}}_{1}|$ in general graphs (i.e., non-trees) as well. The key difference in the general setting is that vertices in $\partial{\mathcal{I}}(t)$ may have multiple infected neighbors. For high-degree vertices, this effect is most pronounced; indeed, as Lemma [ref] shows, it may be the case that a high-degree $v\in\partial{\mathcal{I}}(t)$ has up to $\sqrt{n}$ infected neighbors. Accounting for these qualitative differences from the tree case leads to slight differences in our results for general graphs, as the following result shows.
Lemma 5.2.
Let $G$ be a graph on $n$ vertices, let $\beta>1/2$ and assume that $\gamma>\max\{(1-\beta)/2,1-2\beta\}$. Recall the event ${\mathcal{A}}$ from Definition 4.5. Then for any $t\geq 0$, it holds that
($\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)=o\left(\frac{1}{n}\right).$
Proof.
As in the proof of Lemma [ref], define the event
${\mathcal{A}}_{t}:=\left\{\text{For all $v\in V\setminus{\mathcal{I}}[0,t]$, it holds that $|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq n^{1/2+o(1)}$}\right\}.$
Note that ${\mathcal{A}}_{t}$ is ${\mathcal{F}}_{t}$-measurable, and ${\mathcal{A}}\subset{\mathcal{A}}_{t}$. In particular,
($\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)$ ($\leq\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}_{t}\right|{\mathcal{F}}_{t}\right)$
$=\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\geq n^{\gamma}\right|{\mathcal{F}}_{t}\right)\mathbf{1}({\mathcal{A}}_{t}).$ (5.3)
We therefore proceed by bounding the probability on the right hand side, using the same proof ideas as Lemma 5.1. For each $v\in\partial{\mathcal{I}}(t)$, notice that
$\min_{u\in{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]}F(u,v)\sim\mathrm{Exp}(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|),$
so that
$\operatorname{\mathbb{P}}\left(v\in{\mathcal{I}}_{1}|{\mathcal{F}}_{t}\right)=1-\exp\left(-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\right).$
Moreover, by the independence of edge weights, we have that $|{\mathcal{I}}_{1}|\stackrel{{\scriptstyle d}}{{=}}\sum_{v\in\partial{\mathcal{I}}(t)}X(v)$, where the $X(v)$’s are independent with $X(v)\sim\operatorname{Bern}\left(1-e^{-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|}\right)$. We can then bound
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}|-\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}||{\mathcal{F}}_{t}]\right|\geq\frac{1}{2}n^{\gamma}\right|{\mathcal{F}}_{t}\right)$ $\stackrel{{\scriptstyle(a)}}{{\leq}}2\exp\left(-\frac{n^{2\gamma}/8}{\sum_{v\in\partial{\mathcal{I}}(t)}\left(1-\exp(-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|)\right)+n^{\gamma}/6}\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2\exp\left(-\frac{n^{2\gamma}/8}{n^{1-\beta+o(1)}+n^{\gamma}/6}\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}o\left(\frac{1}{n}\right).$ (5.4)
Above, $(a)$ is due to Bernstein’s inequality. Step $(b)$ uses the inequalities $|\partial{\mathcal{I}}(t)|\leq n$ as well as
$\sum_{v\in\partial{\mathcal{I}}(t)}\left(1-e^{-\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|}\right)\leq\frac{\lambda}{n^{\beta}}\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq\frac{\lambda}{n^{\beta}}\left(dn+nm\right)=n^{1-\beta+o(1)},$ (5.5)
where the second inequality above holds since there are at most $n$ vertices $v$ with $|{\mathcal{N}}(v)|\leq d$ and $m$ vertices $v$ with $|{\mathcal{N}}(v)|\leq n$; the final expression uses that $d=n^{o(1)}$ and $m=n^{o(1)}$. Finally, $(c)$ is due to our assumption that $\gamma>1/2-\beta/2$.
It remains to study $\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}|{\mathcal{F}}_{t}]$. We have that
$\left|\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}||{\mathcal{F}}_{t}]-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|$ $=\left|\sum_{v\in\partial{\mathcal{I}}(t)}\left(1-\exp(\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|)\right)-\frac{\lambda}{n^{\beta}}\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\right|$
$\leq\sum_{v\in\partial{\mathcal{I}}(t)}\left|1-\exp(\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|-\frac{\lambda}{n^{\beta}}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\right|$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\frac{\lambda^{2}}{n^{2\beta}}\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\lambda^{2}n^{1-2\beta+o(1)}$
$\stackrel{{\scriptstyle(e)}}{{\leq}}\frac{1}{2}n^{\gamma}.$ (5.6)
Above, $(c)$ is due to Taylor’s theorem (in particular, the inequality holds as long as $\lambda n^{-\beta}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|=o(1)$, which is the case since ${\mathcal{A}}_{t}$ holds and $\beta>1/2$). Step $(d)$ uses the bound $\sum_{v\in\partial{\mathcal{I}}(t)}|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|\leq n^{1+o(1)}$, which was proved in (5.5). Finally, $(e)$ follows since $\gamma>1-2\beta$.
Together, (5.3), (5.4) and (5.6) prove the lemma. ∎
[next result: bounding $|{\mathcal{I}}_{2}|$ for both trees and general graphs]
Lemma 5.3.
Let $\beta,\gamma>0$ and let $t\geq 0$. If $G$ is a tree, then
$\operatorname{\mathbb{P}}\left(\left.|{\mathcal{I}}_{2}|\geq n^{\gamma}\right|{\mathcal{F}}_{t}\right)\leq n^{-\min\{\beta,\gamma+2\beta-1\}+o(1)}.$
For general $G$, it holds that
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq n^{\gamma}\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)\leq n^{-\min\{\beta-1/2,\gamma+2\beta-1\}+o(1)}.$
Proof.
As a shorthand let $\delta:=n^{-\beta}$, and define the event
${\mathcal{B}}_{t}:=\{\mathsf{highdeg}(G)\cap{\mathcal{I}}(t,t+\delta]=\emptyset\}.$
Observe that for each $u\in{\mathcal{I}}_{2}$, there must exist a distinct path $P$ from ${\mathcal{I}}[0,t]$ to $u$ of length at least 2 and weight at most $\delta$ such that only the first vertex in $P$ is an element of ${\mathcal{I}}[0,t]$. Moreover, if ${\mathcal{B}}_{t}$ holds, then all vertices in $P$ must have degree at most $d$. Otherwise, there would exist a path of weight at most $\delta$ connecting a vertex in ${\mathcal{I}}[0,t]$ to some $u\in\mathsf{highdeg}(G)$, which would imply that $T(u)\in(t,t+\delta]$. However, this cannot happen on the event ${\mathcal{B}}_{t}$.
To see the implications of the discussion above, let ${\mathcal{P}}$ be the set of paths $P$ in $G$ such that $|P|\geq 2$, the first vertex of $P$ is the only one in ${\mathcal{I}}[0,t]$, and all vertices in $P$ have degree at most $d$. Since there exists a distinct path $P\in{\mathcal{P}}$ with weight at most $\delta$ for each $u\in{\mathcal{I}}_{2}$, it holds on the event ${\mathcal{B}}_{t}$ that
$|{\mathcal{I}}_{2}|\leq|\{P\in{\mathcal{P}}:\mathsf{weight}(P)\leq\delta\}|.$
We proceed by bounding the expectation of the right hand side as follows.
$\operatorname{\mathbb{E}}\left[|\{P\in{\mathcal{P}}:\mathsf{weight}(P)\leq\delta\}||{\mathcal{F}}_{t}\right]=\sum_{\ell\geq 2}\sum_{P\in{\mathcal{P}}:|P|=\ell}\operatorname{\mathbb{P}}(\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t})\leq\sum_{\ell\geq 2}|\{P\in{\mathcal{P}}:|P|=\ell\}|(\lambda\delta)^{\ell}$ (5.7)
Above, the final inequality holds since $\mathsf{weight}(P)\sim\mathrm{Gamma}(|P|,\lambda)$. We continue by bounding the number of paths in ${\mathcal{P}}$ of length $\ell$ through elementary combinatorial arguments. If $v\in{\mathcal{I}}[0,t]$ has degree at most $d$, the number of paths $P\in{\mathcal{P}}$ which begin at $v$ and have length $\ell$ is $d^{\ell}$. As there are at most $n$ vertices in ${\mathcal{I}}[0,t]$ of degree at most $d$, the total number of $P\in{\mathcal{P}}$ of length $\ell$ starting at such a vertex is at most $nd^{\ell}$. On the other hand, if $v\in{\mathcal{I}}[0,t]$ is in $\mathsf{highdeg}(G)$, then the total number of paths of length $\ell$ beginning at $v$ is at most $nd^{\ell-1}$, since the degree of $v$ is at most $n$ but all other vertices in the path have degree at most $d$. As there are $m$ vertices in $\mathsf{highdeg}(G)$, the number of such paths is at most $mnd^{\ell-1}$. Combining both cases, we have that
|$|\{P\in{\mathcal{P}}:|P|=\ell\}|\leq nd^{\ell}+mnd^{\ell-1}\leq n^{1+o(1)}d^{\ell}.$
Substituting this bound into (5.7) shows that
$\operatorname{\mathbb{E}}\left[|\{P\in{\mathcal{P}}:\mathsf{weight}(P)\leq\delta\}||{\mathcal{F}}_{t}\right]\leq\sum_{\ell\geq 2}n^{1+o(1)}d^{\ell}(\lambda\delta)^{\ell}=n^{1-2\beta+o(1)},$
where the final expression follows from the geometric sum formula and since $d\lambda\delta=n^{-\beta+o(1)}$. Putting everything together, we have that
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq n^{\gamma}\}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t}\right)\leq\operatorname{\mathbb{P}}\left(|\{P\in{\mathcal{P}}:\mathsf{weight}(P)\leq\delta\}|\geq n^{\gamma}|{\mathcal{F}}_{t}\right)\leq n^{-(\gamma+2\beta-1)+o(1)},$ (5.8)
where the final bound is due to Markov’s inequality. To prove the claim, it remains to study the probability of ${\mathcal{B}}_{t}$. If $G$ is a tree, then
$\operatorname{\mathbb{P}}({\mathcal{B}}_{t}^{c}|{\mathcal{F}}_{t})\leq\sum_{v\in\mathsf{highdeg}(G)\setminus{\mathcal{I}}[0,t]}\operatorname{\mathbb{P}}\left(T(v)\in(t,t+\delta]|{\mathcal{F}}_{t}\right)\leq m\lambda\delta=n^{-\beta+o(1)}.$ (5.9)
Above, the second inequality holds since there are $m$ high-degree vertices and since $T(v)-t$ stochastically dominates a $\mathrm{Exp}(\lambda)$ by Lemma [ref]. Together, (5.8) and (5.9) prove the lemma in the case where $G$ is a tree.
Next, suppose that $G$ is not necessarily a tree. Recalling the event ${\mathcal{A}}$ from Definition 4.5, we have by Lemma [ref] that
$\operatorname{\mathbb{P}}({\mathcal{A}}\cap{\mathcal{B}}_{t}^{c}|{\mathcal{F}}_{t})\leq\sum_{v\in\mathsf{highdeg}(G)}\operatorname{\mathbb{P}}({\mathcal{A}}\cap\{T(v)\in(t,t+\delta]\}|{\mathcal{F}}_{t})\leq mn^{-\left(\beta-\frac{1}{2}\right)+o(1)}=n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
We can therefore bound
($\operatorname{\mathbb{P}}(\{|{\mathcal{I}}_{2}|\geq n^{\gamma}\}\cap{\mathcal{A}}|{\mathcal{F}}_{t})\leq\operatorname{\mathbb{P}}(\{|{\mathcal{I}}_{2}|\geq n^{\gamma}\}\cap{\mathcal{A}}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t})+\operatorname{\mathbb{P}}({\mathcal{A}}\cap{\mathcal{B}}_{t}^{c}|{\mathcal{F}}_{t})]\leq n^{-\min\{\gamma+2\beta-1,\beta-1/2\}+o(1)}.$
[putting everything together]
Lemma 5.4.
Let $v\in V$. Then
$\operatorname{\mathbb{P}}\left(\left||{\mathcal{I}}[T(v),T(v)+n^{-\beta}]|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(v)])|\right|\geq 2n^{\gamma}\right)\leq\begin{cases}n^{-\min\{\beta,\gamma+2\beta-1\}+o(1)},&\text{$G$ is a tree}.\\ n^{-\min\left\{\beta-\frac{1}{2},\gamma+2\beta-1\right\}+o(1)},&G\in...\end{cases}$
Proof.
Fix $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Define the stopping time $T:=T(v)$, as well as the events
${\mathcal{E}}_{t}$ :=$:=\left\{\left||{\mathcal{I}}[t,t+n^{-\beta}]|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|\right|\leq 2n^{\gamma}\right\}$
${\mathcal{E}}$ :=$:=\left\{\left||{\mathcal{I}}[T,T+n^{-\beta}]|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T])|\right|\leq 2n^{\gamma}\right\}.$
Since ${\mathcal{I}}[t,t+n^{-\beta}]={\mathcal{I}}_{1}\cup{\mathcal{I}}_{2}$, Lemmas [ref] imply that if $G$ is a tree,
$\operatorname{\mathbb{P}}({\mathcal{E}}_{t}^{c}|{\mathcal{F}}_{t})\leq n^{-\min\{\beta,\gamma+2\beta-1\}+o(1)}.$
In the general case, we have that
$\operatorname{\mathbb{P}}({\mathcal{E}}_{t}^{c}\cap{\mathcal{A}}|{\mathcal{F}}_{t})\leq n^{-\min\{\beta-1/2,\gamma+2\beta-1\}+o(1)}.$
Since $T$ is almost surely finite, we have that
[…]
if $G$ is a tree. Taking an expectation over ${\mathcal{F}}_{T}$ proves the lemma in the case of trees. Following the same arguments also leads to the claimed result in the general case. ∎
6 Organizing second deriviative stuff
Notation: For $i\in[K]$ and $v\in V$, define
$\widehat{\deg}_{i}(v):=n^{\beta}\left(|{\mathcal{I}}_{i}[T_{i}(v),T_{i}(v)+n^{-\beta}]|-|{\mathcal{I}}_{i}[T_{i}(v)-n^{-\beta},T_{i}(v)]|\right).$
Main result (formulate as algorithm): under some conditions on $\alpha,\beta,K$, it holds with probability $1-o(1)$ that:
  • For $v\in V$ such that $\deg(v)\geq n^{\alpha}$, $\sum_{i=1}^{K}\mathbf{1}(\widehat{\deg}_{i}(v)\geq n^{\alpha})>K/2.$
  • For $v\in V$ such that $\deg(v)\leq d$, $\sum_{i=1}^{K}\mathbf{1}(\widehat{\deg}_{i}(v)\geq n^{\alpha})<K/2.$
[first intermediate result: estimating the cut]
Lemma 6.1.
For any $t\geq 0$, it holds that
($\operatorname{\mathbb{P}}\left(\left\{\left|I\left[t,t+\frac{1}{n^{\beta}}\right]-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}(t))\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-c}$
[estimating changes in the cut]
Lemma 6.2.
Let $\gamma>1-\beta$. Then for any $t\geq 0$,
($\operatorname{\mathbb{P}}\left(\left\{\left|\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{2}}\right)\right)-\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{\beta}}\right)\right)\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-c}$
[discretizing] Let $\mathbb{T}_{n}$ be a discrete set of evenly spaced points that are $1/n^{2}$ apart.
Lemma 6.3.
Let $K\geq...$. With probability $1-o(1)$, it holds for all $t\in\mathbb{T}_{n}$ that
$\sum_{i=1}^{K}\mathbf{1}\left(\left|I\left[t,t+\frac{1}{n^{\beta}}\right]-I\left[t-\frac{1}{n^{\beta}},t\right]-\frac{\lambda}{n^{\beta}}\left(\operatorname{\mathsf{cut}}({\mathcal{I}}(t))-\operatorname{\mathsf{cut}}\left({\mathcal{I}}\left(t-\frac{1}{n^{2}}\right)\right)\right)\right|\leq n^{\gamma}\right)\geq\frac{4K}{5}.$
[from here, we need to relate the stopping times to the discretization. The result should then follow immediately.]
[good notation for empirical second derivative?]
[non-tree results below]
[first main result: computing the cut]
Lemma 6.4.
Let $v\in V$ and $\beta>1/2$. It holds that
($\operatorname{\mathbb{P}}\left(\left\{\left|\left|{\mathcal{I}}\left(T(v),T(v)+\frac{1}{n^{\beta}}\right]\right|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(v)])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right)\leq n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
[From here on out, the proof is pretty easy. If you look at $\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(v)])-\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(u)]$ for consecutively infected vertices $v,u$, you’ll reveal the degree info. After all, if vertices are consecutive, you extract only the effect of the added vertex.]
6.1 Analysis of $|{\mathcal{I}}_{1}|$: Proof of Lemmas 5.1 and 6.4
Our next result shows that with high probability, ${\mathcal{I}}_{2}=\emptyset$.
Lemma 6.5.
Suppose that $\beta>1/2$. Then for any $t\geq 0$, it holds that
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)\leq n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
Proof.
Let $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Observe that if ${\mathcal{I}}_{2}\neq\emptyset$, then there must exist a path $P$ of length 2 and weight at most $\delta$ for which only the first vertex in $P$ is an element of ${\mathcal{I}}[0,t]$. Additionally, if we define the event
${\mathcal{B}}_{t}:=\{\mathsf{highdeg}(G)\cap{\mathcal{I}}(t,t+\delta]=\emptyset\},$
then all vertices in $P$ must have degree at most $d$. Otherwise, there would exist a path of weight at most $\delta$ connecting a vertex in ${\mathcal{I}}[0,t]$ to some $u\in\mathsf{highdeg}(G)$, which would imply that $T(u)\in(t,t+\delta]$. However, this cannot happen on the event ${\mathcal{B}}_{t}$. Importantly, ${\mathcal{B}}_{t}$ holds with high probability provided ${\mathcal{A}}$ holds; indeed, we have that
$\operatorname{\mathbb{P}}({\mathcal{A}}\cap{\mathcal{B}}_{t}^{c}|{\mathcal{F}}_{t})\leq\sum_{v\in\mathsf{highdeg}(G)}\operatorname{\mathbb{P}}({\mathcal{A}}\cap\{T(v)\in(t,t+\delta]\}|{\mathcal{F}}_{t})\leq mn^{-\left(\beta-\frac{1}{2}\right)+o(1)}=n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
To see the implications of the discussion above, let ${\mathcal{Q}}$ be the set of paths $P$ in $G$ of length 2 for which the first vertex in $P$ is the only one in ${\mathcal{I}}[0,t]$, and all vertices in $P$ have degree at most $d$.
Notice that we can bound
$|{\mathcal{Q}}|\leq nd^{2}+mnd\leq 2mnd^{2}.$ (6.1)
The first inequality can be explained as follows. If $v\in{\mathcal{I}}[0,t]$ and $v$ has degree at most $d$, the number of paths $P\in{\mathcal{Q}}$ which begin at $v$ is at most $d^{2}$. As there are at most $n$ vertices in ${\mathcal{I}}[0,t]$ of degree at most $d$, the total number of $P\in{\mathcal{Q}}$ starting at such a vertex is $nd^{2}$. On the other hand, if $v\in\mathsf{highdeg}(G)\cap{\mathcal{I}}[0,t]$, then the total number of paths of length 2 beginning at $v$ is at most $nd$, since $\deg(v)\leq n$. As there are $m$ vertices in $\mathsf{highdeg}(G)$, the number of paths in ${\mathcal{Q}}$ for this case is at most $mnd$. Summing both together leads to the claimed bound.
By our arguments above,
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t}\right)$ $\leq\operatorname{\mathbb{P}}\left(\exists P\in{\mathcal{Q}}:\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t}\right)$
$\leq\sum_{P\in{\mathcal{Q}}}\operatorname{\mathbb{P}}(\mathsf{weight}(P)\leq\delta|{\mathcal{F}}_{t})$
$\stackrel{{\scriptstyle(a)}}{{\leq}}|{\mathcal{Q}}|\lambda^{2}\delta^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2mnd^{2}\lambda^{2}\delta^{2}=n^{-(2\beta-1)+o(1)}.$
Above, $(a)$ follows since $\mathsf{weight}(P)\sim\mathrm{Gamma}(2,\lambda)$ and $(b)$ uses (6.1).
Putting everything together, we have that
($\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)$ ($\leq\operatorname{\mathbb{P}}\left(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}\cap{\mathcal{B}}_{t}|{\mathcal{F}}_{t}\right)+\operatorname{\mathbb{P}}({\mathcal{A}}\cap{\mathcal{B}}_{t}^{c}|{\mathcal{F}}_{t})$
$\leq n^{-(2\beta-1)+o(1)}+n^{-\left(\beta-1/2\right)+o(1)}=n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
We now put everything together to prove the desired result.
Proof of Lemma 6.4 .
For any $t\geq 0$, Lemmas 5.2 and 6.5 imply that
($\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}(t,t+\delta]-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)\\ \leq\operatorname{\mathbb{P}}\left(\left.\left\{\left||{\mathcal{I}}_{1}|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{t}\right)+\operatorname{\mathbb{P}}(\{|{\mathcal{I}}_{2}|\geq 1\}\cap{\mathcal{A}}|{\mathcal{F}}_{t})\leq n^{-\left(\beta-\frac{1}{2}\right)+o(1)}.$
Through identical reasoning as the proof of Lemma [reF], it follows that
($\operatorname{\mathbb{P}}\left(\left.\left\{\left|\left|{\mathcal{I}}\left(T(v),T(v)+\frac{1}{n^{\beta}}\right]\right|-\frac{\lambda}{n^{\beta}}\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])\right|\geq n^{\gamma}\right\}\cap{\mathcal{A}}\right|{\mathcal{F}}_{T(v)}\right)\leq n^{-\left(\beta-\frac{1}{2}\right)+o(1)},$
which holds since $T(v)<\infty$ almost surely. Finally, the desired result follows from taking an expectation with respect to ${\mathcal{F}}_{T(v)}$. ∎
7 Algorithm that works for all $\alpha\in(0,1)$
[high-level idea] For each vertex, look at the number of consistent infections that occur in 1 timestep after it is infected. If we look at the set of vertices that have at least $D/2$ vertices in their consistent infection set, this will be a subgraph of size $n^{\alpha+\epsilon}$. We can then do the same second-derivative test to find the super-spreaders.
[update: this doesn’t work!!] At least not for general graphs, due to the many possible paths that can lead to the shortest one. So let’s just throw in the proof for trees, and emphasize the difficulties. General graphs can be an open problem.
[formal results]
[a few results on comparing gamma variables]
[possible notation:] When there is a single high-degree vertex, denote it by $v_{\mathrm{HD}}$.
Lemma 7.1.
Suppose that $X\sim\mathrm{Gamma}(a,\lambda)$ and $Y\sim\mathrm{Gamma}(b,\lambda)$ are independent with $1001\leq a\leq b/$ 100. Then
$\operatorname{\mathbb{P}}(Y\leq X-1)\leq e^{\lambda}\left(\frac{2}{3}\right)^{a+b}.$
Proof.
For any $\theta\in(0,\lambda)$, it holds that
$\operatorname{\mathbb{P}}(Y\leq X-1)=\operatorname{\mathbb{P}}(X-Y\geq-1)\leq e^{\theta}\operatorname{\mathbb{E}}[e^{\theta(X-Y)}]\leq e^{\lambda}\left(1-\frac{\theta}{\lambda}\right)^{-a}\left(1+\frac{\theta}{\lambda}\right)^{-b}$
Setting $\theta:=\lambda(b-a)/(a+b)$, we arrive at the inequality
$\operatorname{\mathbb{P}}(Y\leq X-1)\leq\exp\left\{\lambda-(a+b)\left(\log 2-H\left(\frac{a}{a+b}\right)\right)\right\},$ (7.1)
where we define the binary entropy function $H(p):=-p\log p-(1-p)\log(1-p)$. Since $b/a\geq 100$, we have that
$H\left(\frac{a}{a+b}\right)\leq H\left(\frac{1}{101}\right)\leq\log\left(\frac{4}{3}\right),$
and it follows that
$\log 2-H\left(\frac{a}{a+b}\right)\geq\log\left(\frac{3}{2}\right).$
Substituting this inequality into (7.1) implies the claimed result. ∎
Lemma 7.2.
Let $X\sim\mathrm{Gamma}(a,\lambda)$ with $a\geq 1$. For $a$ sufficiently large, it holds that
$\sup_{x>0}\operatorname{\mathbb{P}}\left(|X-x|\leq 1\right)\leq\frac{\lambda}{\sqrt{a}}.$
Proof.
Letting $f(x)$ be the density function for $X$, define $f_{\sup}:=\sup_{x>0}f(x)$ to be its maximum value. Since we can bound
[…]
it suffices to characterize $f_{\sup}$. To this end, we have that
$f_{\sup}$ $:=\sup_{x>0}\frac{\lambda^{a}}{\Gamma(a)}x^{a-1}e^{-\lambda x}\stackrel{{\scriptstyle(a)}}{{=}}\frac{\lambda}{\Gamma(a)}\left(a-1\right)^{a-1}e^{-(a-1)}\stackrel{{\scriptstyle(b)}}{{\sim}}\frac{\lambda}{\sqrt{2\pi a}}.$
Above, $(a)$ uses the fact that the mode of a $\mathrm{Gamma}(a,\lambda)$ random variable is $(a-1)/\lambda$, and $(b)$ is due to Stirling’s approximation. Substituting this asymptotic expression into (7.2) shows that for $a$ sufficiently large,
$\operatorname{\mathbb{P}}(|X-x|\leq 1)\lesssim\frac{2\lambda}{\sqrt{2\pi a}}\leq\frac{\lambda}{\sqrt{a}}.$
[narrowing down the search space]
[need to set the right value of $R$ and $K$]
Lemma 7.3.
It holds with probability $1-o(1)$ that for all $v\in V$,
$\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\subset{\mathcal{N}}(v,R).$
Proof.
As a shorthand, we set $R:=\epsilon\log(n)/\log(d)$ in this proof. Fix vertices $u,v\in V$ with $\mathsf{dist}_{G}(u,v)\geq R$ as well as $i\in[K]$. Let $P_{i}$ and $Q_{i}$ be the unique paths from the source vertex $s_{i}$ to $u_{i}$ and $v_{i}$, respectively. If $P_{i}$ and $Q_{i}$ are disjoint, then it holds that $T_{i}(u)=\mathsf{weight}_{i}(P_{i})$ and $T_{i}(v)=\mathsf{weight}_{i}(Q_{i})$ are independent, with $\mathsf{weight}_{i}(P_{i})\sim\mathrm{Gamma}(|P_{i}|,\lambda)$ and $\mathsf{weight}_{i}(Q_{i})\sim\mathrm{Gamma}(|Q_{i}|,\lambda)$. As the graph is a tree, it holds that $|P_{i}|+|Q_{i}|=\mathsf{dist}_{G}(u,v)\geq R$, hence we have that either $|P_{i}|\geq R/2$ or $|Q_{i}|\geq R/2$. Lemma 7.2 therefore implies
$\operatorname{\mathbb{P}}\left(|T_{i}(u)-T_{i}(v)|\leq 1\right)\leq\frac{\lambda}{\sqrt{R/2}}.$ (7.3)
In the case where $P_{i}$ and $Q_{i}$ are not disjoint, let $w$ be the last common element of both paths. We then define $P_{i}^{\prime}$ and $Q_{i}^{\prime}$ to be the portion of $P_{i}$ and $Q_{i}$, respectively, that starts at $w$ instead of $s_{i}$. By definition, $P_{i}^{\prime}$ and $Q_{i}^{\prime}$ are disjoint, and we have that $T_{i}(u)=T_{i}(w)+\mathsf{weight}_{i}(P_{i}^{\prime})$ and $T_{i}(v)=T_{i}(w)+\mathsf{weight}_{i}(Q_{i}^{\prime})$. Moreover, $\mathsf{weight}_{i}(P_{i}^{\prime})\sim\mathrm{Gamma}(|P_{i}^{\prime}|,\lambda)$ and $\mathsf{weight}(Q_{i}^{\prime})\sim\mathrm{Gamma}(|Q_{i}^{\prime}|,\lambda)$, are independent, and as before, $|P_{i}^{\prime}|+|Q_{i}^{\prime}|=\mathsf{dist}_{G}(u,v)\geq R$. By the same logic as the case where $P_{i},Q_{i}$ are disjoint, we see that (7.3) holds in this case as well.
From (7.3), it is readily seen that for $K$ satisfying [eqref], it holds that
$\operatorname{\mathbb{P}}\left(|T_{i}(u)-T_{i}(v)|\leq 1,\forall i\in[K]\right)=o\left(\frac{1}{n^{2}}\right).$ (7.4)
Finally, to prove the claim, we can bound
$\operatorname{\mathbb{P}}\left(\exists v\in V:\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\not\subset{\mathcal{N}}(v,R)\right)$ $=\operatorname{\mathbb{P}}\left(\sum_{v\in V}\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\setminus{\mathcal{N}}(v,R)\right|\geq 1\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{v\in V}\operatorname{\mathbb{E}}\left[\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\setminus{\mathcal{N}}(v,R)\right|\right]$
$=\sum_{v\in V}\sum_{u\in V\setminus{\mathcal{N}}(v,R)}\operatorname{\mathbb{P}}\left(u\in\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\right)$
$\leq\sum_{v\in V}\sum_{u\in V\setminus{\mathcal{N}}(v,R)}\operatorname{\mathbb{P}}\left(\left|T_{i}(u)-T_{i}(v)\right|\leq 1,\forall i\in[K]\right)$
=$\stackrel{{\scriptstyle(b)}}{{=}}o(1).$
Above, $(a)$ is due to Markov’s inequality, and $(b)$ follows from (7.4). ∎
Define the set
$\widehat{U}:=\left\{v\in V:\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\right|>n^{\alpha/2}\right\}.$
[set $R$ below]
Lemma 7.4.
With probability $1-o(1)$, the following hold:
  1. 1.
    $\widehat{U}\subset{\mathcal{N}}(v_{\mathrm{HD}},R)$.
  2. 2.
    $v_{\mathrm{HD}}\in\widehat{U}$.
  3. 3.
    $|\widehat{U}\cap{\mathcal{N}}(v_{\mathrm{HD}})|\geq\frac{1}{4}(1-e^{-\lambda/2})^{K}|{\mathcal{N}}(v_{\mathrm{HD}})|$.
Proof.
Suppose that $u\notin{\mathcal{N}}(v_{\mathrm{HD}},R)$, which is equivalent to $v_{\mathrm{HD}}\notin{\mathcal{N}}(u,R)$. By Lemma 7.3, it holds with probability $1-o(1)$ that for all such $u$,
$\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(u),T_{i}(u)+1]\right|\leq|{\mathcal{N}}(u,R)|\leq d^{R}=o(n^{\alpha/2}).$ (7.5)
[explain upper bound on $|{\mathcal{N}}(u,R)|$] In particular, (7.5) implies that $\widehat{U}$ is contained in ${\mathcal{N}}(v_{\mathrm{HD}},R)$ with probability $1-o(1)$, proving Item #1.
Next, we turn to the proof of Item #2. For each $i\in[K]$, we let $v_{i}$ be the penultimate vertex on the unique path from $s_{i}$ to $v_{\mathrm{HD}}$ if $s_{i}\neq v$, otherwise we set $v_{i}:=v_{\mathrm{HD}}$. Notice that for $u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{i}\}$, it holds that
$T_{i}(u)=T_{i}(v_{\mathrm{HD}})+E_{i}(u),\qquad u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{i}\},i\in[K],$
where the $E_{i}(u)$’s are i.i.d. $\mathrm{Exp}(\lambda)$ random variables. Notice that for $u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}$,
$E_{i}(u)\leq 1,\forall i\in[K]\Rightarrow u\in\bigcap_{i=1}^{K}{\mathcal{I}}_{i}[T_{i}(v_{\mathrm{HD}}),T_{i}(v_{\mathrm{HD}})+1].$
Since $\operatorname{\mathbb{P}}(\max_{i\in[K]}E_{i}(u)\leq 1)=(1-e^{-\lambda})^{K}$, the $E_{i}(u)$’s are i.i.d., and $|{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}|\geq n^{\alpha}-K\geq n^{\alpha}/2$, it follows that $|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]|$ stochastically bounds a $\mathrm{Bin}(n^{\alpha}/2,(1-e^{-\lambda})^{K})$ random variable. It now follows from Bernstein’s inequality that
$\operatorname{\mathbb{P}}\left(\left|\bigcap_{i=1}^{K}{\mathcal{I}}[T_{i}(v),T_{i}(v)+1]\right|\leq\frac{1}{4}(1-e^{-\lambda})^{K}n^{\alpha}\right)=o(1).$
Item #2 immediately follows, since $(1-e^{-\lambda})^{K}n^{\alpha}/4\geq n^{\alpha/2}$ for $n$ sufficiently large.
We now prove Item #[ref]. Let us define the sets
$S_{1}$ :=$:=\{u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}:E_{i}(u)\in[0,1/2],\forall i\in[K]\}$
$S_{2}$ :=$:=\{u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}:E_{i}(u)\in[1/2,1],\forall i\in[K]\}.$
For $u\in S_{1}$ and $w\in S_{2}$, it holds for all $i\in[K]$ that $T_{i}(w)\in[T_{i}(u),T_{i}(u)+1]$, hence
$S_{2}\subset\bigcap_{i=1}^{K}{\mathcal{I}}_{i}[T_{i}(u),T_{i}(u)+1],\qquad,\forall u\in S_{1}.$
In particular, [eqref] implies that $|\widehat{U}\cap{\mathcal{N}}(v_{\mathrm{HD}})|\geq|S_{1}|$ if $|S_{2}|\geq n^{\alpha/2}$. We therefore proceed by establishing lower bounds for $|S_{1}|$ and $|S_{2}|$. For $u\in{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}$, it holds that
$\operatorname{\mathbb{P}}(u\in S_{1})$ $=\prod_{i=1}^{K}\operatorname{\mathbb{P}}(E_{i}(u)\in[0,1/2])=(1-e^{-\lambda/2})^{K}$
$\operatorname{\mathbb{P}}(u\in S_{2})$ $=\prod_{i=1}^{K}\operatorname{\mathbb{P}}(E_{i}(u)\in[1/2,1])=\left(e^{-\lambda/2}(1-e^{-\lambda/2})\right)^{K}.$
As the $E_{i}(u)$’s are independent, we may apply Bernstein’s inequality to obtain that
$|S_{1}|$ $\geq\frac{1}{2}(1-e^{-\lambda/2})^{K}|{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}|\geq\frac{1}{4}(1-e^{-\lambda/2})^{K}|{\mathcal{N}}(v_{\mathrm{HD}})|$
$|S_{2}|$ $\geq\frac{1}{2}\left(e^{-\lambda/2}(1-e^{-\lambda/2})\right)^{K}|{\mathcal{N}}(v_{\mathrm{HD}})\setminus\{v_{1},\ldots,v_{K}\}|\geq\frac{1}{4}\left(e^{-\lambda/2}(1-e^{-\lambda/2})\right)^{K}|{\mathcal{N}}(v_{\mathrm{HD}})|,$
with probability $1-o(1)$. In the final inequality on both lines, we have used that |$|\{v_{1},\ldots,v_{K}\}|\leq K\leq|{\mathcal{N}}(v_{\mathrm{HD}})|/2$. In particular, $|S_{2}|\geq n^{\alpha/2}$ with probability $1-o(1)$, so it follows that
$|\widehat{U}\cap{\mathcal{N}}(v_{\mathrm{HD}})|\geq|S_{1}|\geq\frac{1}{4}(1-e^{-\lambda/2})^{K}|{\mathcal{N}}(v_{\mathrm{HD}})|,$
with probability $1-o(1)$. ∎
7.1 Proof of restricted cut thing
[next result: for a given set $S$, $|{\mathcal{I}}[T(v),T(v)+\delta]\cap S|$ concentrates around $|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,T(v)]\cap S)|$.]
Let $S\subset V$, fix $t\geq 0$, and condition on ${\mathcal{F}}_{t}$. Recall the definitions of ${\mathcal{I}}_{1}$ and ${\mathcal{I}}_{2}$ from Section [ref] (see [eqrefs]); the objective of this section is to study $|{\mathcal{I}}_{1}\cap S|$ and $|{\mathcal{I}}_{2}\cap S|$. The techniques needed for doing so are largely the same as those in Section [ref], and we emphasize the key differences here.
We begin by studying the size of ${\mathcal{I}}_{1}\cap S$.
Lemma 7.5.
Let $\gamma>\max\{\alpha/2,1-2\beta\}$ and let $t\geq 0$. Then it holds with probability $1-o(1/n)$ that
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}\cap\widehat{U}|-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|\right|\geq n^{\gamma}\right|{\mathcal{F}}_{t}\right)=o\left(\frac{1}{n}\right).$
Remark 7.6.
The event ${\mathcal{A}}$ is not needed here, since in trees, there is a single infection path from the source to any given vertex.
Proof.
Condition on ${\mathcal{F}}_{t}$ as well as $\widehat{U}$. Since $G$ is a tree, we have for each $v\in\partial{\mathcal{I}}(t)$ that $v$ has exactly one infected neighbor at time $t$. Hence $T(v)-t\sim\mathrm{Exp}(\lambda)$ conditioned on ${\mathcal{F}}_{t}$, so
$\operatorname{\mathbb{P}}(v\in{\mathcal{I}}_{1}|{\mathcal{F}}_{t})=\operatorname{\mathbb{P}}(T(v)-t\in[0,n^{-\beta}]|{\mathcal{F}}_{t})=1-e^{-\lambda/n^{\beta}}.$
By the conditional independence of infection times in $\partial{\mathcal{I}}(t)$ with respect to ${\mathcal{F}}_{t}$ and $\widehat{U}$, we have that $|{\mathcal{I}}_{1}|\sim\operatorname{Bin}(|\partial{\mathcal{I}}(t)|,1-e^{-\lambda/n^{\beta}})$.
By the definition of ${\mathcal{I}}_{1}$, we have that ${\mathcal{I}}_{1}\cap S\subset\partial{\mathcal{I}}(t)\cap\{v\in V:{\mathcal{N}}(v)\cap\widehat{U}\neq\emptyset\}$. Observe further that each $v\in\partial{\mathcal{I}}(t)$ has exactly one neighbor in ${\mathcal{I}}[0,t]$. Indeed, if there were at least two infected neighbors of $v$, there must exist two distinct paths from the cascade source $s$ to $v$, which is not possible under our assumption that the underlying graph is a tree. Hence it holds for $v\in\partial{\mathcal{I}}(t)$ that
$\operatorname{\mathbb{P}}(v\in{\mathcal{I}}_{1}|{\mathcal{F}}_{t})=1-e^{-\lambda\delta}.$
Following the proof of Lemma [ref], we have the representation
$|{\mathcal{I}}_{1}\cap\widehat{U}|\stackrel{{\scriptstyle d}}{{=}}\sum_{v\in\partial{\mathcal{I}}(t):{\mathcal{N}}(v)\cap\widehat{U}\neq\emptyset}X(v),$
where the $X(v)$’s are independent with $X(v)\sim\mathrm{Bern}\left(1-e^{-\lambda\delta}\right)$. Analogously to [eqref], Bernstein’s inequality implies that for any $x\geq 0$,
$\operatorname{\mathbb{P}}\left(\left.\left||{\mathcal{I}}_{1}\cap\widehat{U}|-\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}\cap\widehat{U}||{\mathcal{F}}_{t},\widehat{U}]\right|\geq x\right|{\mathcal{F}}_{t},\widehat{U}\right)\leq 2\exp\left(-\frac{-x^{2}/2}{|\partial{\mathcal{I}}(t)\cap\{v:{\mathcal{N}}(v)\cap\widehat{U}\neq\emptyset\}|\left(1-e^{-\lambda\delta}\right)+x/3}\right).$
To simplify the bound further, we make a number of observations. First, since the graph $G$ is a tree, we have that
$|\partial{\mathcal{I}}(t)\cap\{v\in V:{\mathcal{N}}(v)\cap\widehat{U}\neq\emptyset\}|=|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|.$
Moreover, we can bound
$|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|\leq|\operatorname{\mathsf{cut}}(\widehat{U})|\leq\sum_{v\in\widehat{U}}|{\mathcal{N}}(v)|\leq|{\mathcal{N}}(v_{\mathrm{HD}})|+d|\widehat{U}|\leq n^{\alpha+\epsilon}.$
[explain.] As a consequence of [eqref], we can characterize the conditional expectation of $|{\mathcal{I}}_{1}\cap\widehat{U}|$ as follows.
$\left|\operatorname{\mathbb{E}}[|{\mathcal{I}}_{1}\cap\widehat{U}||{\mathcal{F}}_{t},\widehat{U}]-\frac{\lambda}{n^{\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|\right|\leq\frac{\lambda^{2}}{n^{2\beta}}|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t]\cap\widehat{U})|\leq\lambda^{2}n^{\alpha-2\beta+\epsilon}\leq\frac{1}{2}n^{\gamma}.$
[explain] Setting $x=n^{\gamma}/2$ in [eqref] as well as using the inequalities [eqref] and $1-e^{-\lambda\delta}\leq\lambda\delta$ proves the claim. ∎
[Next result: bounding $|{\mathcal{I}}_{2}\cap\widehat{U}|$]
8 Lower bounds for $\alpha\in(0,1/2)$
Theorem 8.1.
Fix $\epsilon>0$ and $\alpha\in(0,1/2)$. Let $D=n^{\alpha+o(1)}$ and let ${\mathcal{G}}$ be the class of all graphs where all vertices have degree at most $\log^{2}n$, except for at most one vertex with degree at least $D$. If
$K\leq\left(\frac{1-2\alpha-\epsilon}{5}\right)\frac{\log n}{\log\log n},$
then
$\inf_{\widehat{v}}\max_{G\in{\mathcal{G}}}\operatorname{\mathbb{P}}_{G}\left(\widehat{v}(\mathbf{T})\neq\mathsf{highdeg}(G)\right)=1-o(1).$
To prove the theorem, we will construct a tailored family of graphs ${\mathcal{H}}\subset{\mathcal{G}}$ for which we can tractably analyze the distribution of infection times, while also constituting a “hard" instance for estimating high-degree vertices. To construct ${\mathcal{H}}$, we first describe a particular deterministic graph $H$; see the definition below as well as Figure [todo] for a visualization.
Definition 8.2.
Fix a positive integer $N$. Let $H^{\prime}$ be any connected graph with $N$ vertices, maximum degree at most 3, and diameter at most $2\log N$ (e.g., a balanced binary tree). For each vertex in $u\in H^{\prime}$, we add a path of length $\log^{5}n$ starting at $u$; all added paths are disjoint. We call the resulting graph $H$. The set of endpoints of the paths added to $H^{\prime}$, which are also the set of leaves of $H$, is denoted by $L$.
The addition of paths which are much longer than the diameter of $H^{\prime}$ is a crucial aspect of our construction. At a high level, it ensures that the infection times of vertices in $L$ are primarily determined by the infection times of vertices in the corresponding path, rather than the infection times of vertices in $H^{\prime}$. Since all the added paths are disjoint, the infection times of vertices in $L$ become almost independent, which simplifies much of our analysis. A bit more formally, we have the following useful observation.
Fact 8.3.
The infection times of vertices in $L$, given by the collection $\{\mathbf{T}(v)\}_{v\in L}$, are conditionally independent with respect to $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$, the collection of infection times of vertices in $H^{\prime}$.
To construct the graphs of interest, we will introduce additional vertices to $H$ which connect to $L$ in a particular manner. See the formal definition below as well as the visualization in Figure [todo].
Definition 8.4.
Suppose that $N$ vertices, labelled from 1 to $N$, are added to $H$. Each vertex in $[N]$ forms an edge with a single element of $L$. Define ${\mathcal{H}}_{\emptyset}$ to be the set of resulting graphs for which all vertices in $L$ have degree at most $\log^{2}n$. For each $v\in L$, define ${\mathcal{H}}_{v}$ to be the set of graphs for which each $u\in L\setminus\{v\}$ has degree at most $\log^{2}n$, but $v$ has degree at least $D/2$. The union of all such graphs is given by ${\mathcal{H}}:=\bigcup_{x\in L\cup\{\emptyset\}}{\mathcal{H}}_{x}$.
We make a few simple observations about the family ${\mathcal{H}}$.
  • If $G\in{\mathcal{H}}_{\emptyset}$ then $\mathsf{highdeg}(G)=\emptyset$, and if $G\in{\mathcal{H}}_{v}$ for $v\in L$ then $\mathsf{highdeg}(G)=v$.
  • The total number of vertices in $G\in{\mathcal{H}}$ is $n:=2N+N\log^{5}N$. In particular, $N=n^{1-o(1)}$.
The first step in our analysis is to show that it is challenging to even detect whether the set of observed infection times are due to a graph in ${\mathcal{H}}_{\emptyset}$ or in ${\mathcal{H}}_{v}$ for some $v\in L$. To do so, we construct convenient distributions $\mu_{\emptyset}$ over elements of ${\mathcal{H}}_{\emptyset}$ and $\mu_{v}$ over elements of ${\mathcal{H}}_{v},v\in L$. We let $P_{\emptyset}$ and $P_{v},v\in L$ denote the induced distributions over the infection times. That is, for $x\in L\cup\{\emptyset\}$, $P_{x}$ is the distribution corresponding to $\mathcal{SI}(G,v_{0})^{\otimes K}$ where $G\sim\mu_{x}$. The precise description of the distributions $\{\mu_{x}\}_{x\in L\cup\{\emptyset\}}$ is somewhat complicated, so we defer the details to Section 8.1.
[play up the natural connections to the sparse mixture problem!]
Lemma 8.5.
Let $\epsilon>0$ and $\alpha\in(0,1/2)$. Suppose that $D\leq n^{\alpha}$ for some $\alpha\in(0,1/2)$ and that
$K\leq\left(\frac{1-2\alpha-\epsilon}{5}\right)\frac{\log n}{\log\log n}.$
Then it holds for all $v\in L$ that $\mathrm{TV}(P_{\emptyset},P_{v})=o(1)$.
At a high level, we prove Lemma 8.5 by studying the $\chi^{2}$-divergence between $P_{\emptyset}$ and $P_{v}$ for $v\in L$, which provides an upper bound for the total variation distance and has a useful tensorization property which we leverage in our analysis. For full details, see Section 8.1.
With Lemma 8.5 in hand, we can now prove the theorem.
Proof of Theorem 8.1.
Denote $\delta:=\max_{u,v\in L\cup\{\emptyset\}}\mathrm{TV}(Q_{u},Q_{v})$, and let $\hat{v}$ be any estimator. Then for any $u\in B\cup\{0\}$,
$P_{u}(\hat{v}=u)$ $=1-\sum_{v\in L\cup\{\emptyset\}:v\neq u}P_{u}(\hat{v}=v)$
$\leq 1-\sum_{v\in L\cup\{\emptyset\}:v\neq u}\left(P_{v}(\hat{v}=v)-\delta\right)$
$=1+\delta N-N\min_{v\in L\cup\{\emptyset\}}P_{v}(\hat{v}=v).$
As the upper bound holds for all $u\in L\cup\{\emptyset\}$, it follows that
$\min_{v\in L\cup\{\emptyset\}}P_{v}(\hat{v}=v)\leq 1+\delta N-N\min_{v\in L\cup\{\emptyset\}}P_{v}(\hat{v}=v).$
Rearranging terms shows that
$\min_{v\in L\cup\{\emptyset\}}P_{v}(\hat{v}=v)\leq\frac{1+\delta N}{1+N}\leq\frac{1}{N}+\delta=o(1),$ (8.1)
where the final inequality holds for $n$ sufficiently large in light of Lemma 8.5. Importantly, (8.1) implies the existence of $G\in{\mathcal{H}}$ for which $\operatorname{\mathbb{P}}_{G}(\hat{v}=\mathsf{highdeg}(G))\leq\epsilon$. As this conclusion holds for any estimator $\widehat{v}$, it holds that
$\sup_{\widehat{v}}\min_{G\in{\mathcal{G}}}\operatorname{\mathbb{P}}_{G}(\widehat{v}=v)=o(1),$
which is equivalent to the desired claim. ∎
8.1 Impossibility of detecting super-spreaders
8.1.1 Distributions of graphs in ${\mathcal{H}}$
Before proving the lemma, we must describe the distributions $\{\mu_{x}\}_{x\in L\cup\{\emptyset\}}$ over graphs in ${\mathcal{H}}$. To this end, it is useful to first define corresponding “auxiliary" distributions $\{\mu_{x}^{\prime}\}_{x\in L\cup\{\emptyset\}}$. We say that $G\sim\mu_{\emptyset}^{\prime}$ if each vertex in $[N]$ chooses a uniform random element of $L$ with which they form an edge. On the other hand, we say that $G\sim\mu_{v}^{\prime}$ if, for each $u\in[N]$, $u$ connects to a uniform random element of $L$ with probability $1-D/N$, otherwise $u$ connects to $v$ with probability $D/N$. Finally, for $x\in L\cup\{\emptyset\}$, the measure $\mu_{x}$ is equal to $\mu_{x}^{\prime}$ conditioned on the sampled graph being an element of ${\mathcal{H}}_{x}$.
For $x\in L\cup\{\emptyset\}$, we recall that $P_{x}$ is the distribution corresponding to $\mathcal{SI}(G,v_{0})^{\otimes K}$, where $G\sim\mu_{x}$. Similarly, we define $P_{x}^{\prime}$ to be the distribution corresponding to $\mathcal{SI}(G,v_{0})^{\otimes K}$, where $G\sim\mu_{x}^{\prime}$. The following result shows that the total variation distance between $P_{x}$ and $P_{x}^{\prime}$ is small.
Lemma 8.6.
It holds for any $x\in L\cup\{\emptyset\}$ that $\mathrm{TV}(P_{x},P_{x}^{\prime})=o(1)$.
The lemma implies that we can essentially replace $P_{x}$ with $P_{x}^{\prime}$ in our analysis moving forward. Indeed, by the triangle inequality for the total variation distance, we have for any $v\in L$ that
$\mathrm{TV}(P_{\emptyset},P_{v})\leq\mathrm{TV}(P_{\emptyset}^{\prime},P_{v}^{\prime})+\mathrm{TV}(P_{\emptyset}^{\prime},P_{\emptyset})+\mathrm{TV}(P_{v}^{\prime},P_{v})=\mathrm{TV}(P_{\emptyset}^{\prime},P_{v}^{\prime})+o(1),$ (8.2)
where the final equality is due to Lemma 8.6. Hence it suffices to bound the total variation between the auxiliary distributions.
Proof of Lemma 8.6.
Let $x\in L\cup\{\emptyset\}$. By the data processing inequality for the total variation distance, it holds that $\mathrm{TV}(P_{x},P_{x}^{\prime})\leq\mathrm{TV}(\mu_{x},\mu_{x}^{\prime})$. By the definition of $\mu_{x}$ and $\mu_{x}^{\prime}$, the two distributions can be coupled precisely when $G\sim\mu_{x}^{\prime}$ satisfies $G\in{\mathcal{H}}_{x}$. Hence
$\mathrm{TV}(P_{x},P_{x}^{\prime})\leq\mathrm{TV}(\mu_{x},\mu_{x}^{\prime})=\mu_{x}^{\prime}({\mathcal{H}}_{x}^{c}).$ (8.3)
In light of (8.3), we proceed by proving that $\mu_{x}^{\prime}({\mathcal{H}}_{x}^{c})=o(1)$. Notice that for $G\sim\mu_{x}^{\prime}$, all vertices except for those in $L$ have a deterministic degree which is at most 3, hence it suffices to study the degrees of vertices in $L$ alone. In the case $x=\emptyset$, it holds for every $v\in L$ that $\deg_{G}(v)-1\sim\operatorname{Bin}(N,1/N)$. Bernstein’s inequality implies
$\mu_{\emptyset}^{\prime}\left(\deg_{G}(v)\geq\log^{2}N\right)\leq\mu_{\emptyset}^{\prime}\left(\deg_{G}(v)-1\geq 1+\frac{1}{2}\log^{2}N\right)\leq\mathrm{exp}\left(-\frac{2}{3}\log^{2}N\right).$ (8.4)
It follows that
$\mu_{\emptyset}^{\prime}\left({\mathcal{H}}_{\emptyset}^{c}\right)$ $=\mu_{\emptyset}^{\prime}\left(\exists v\in L:\deg_{G}(v)>\log^{2}n\right)$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\mu_{\emptyset}^{\prime}\left(\exists v\in L:\deg_{G}(v)>\log^{2}N\right)$
$\stackrel{{\scriptstyle(b)}}{{\leq}}N\exp\left(-\frac{2}{3}\log^{2}N\right)=o(1).$ (8.5)
Above, $(a)$ follows since $N\leq n$, and $(b)$ is due to a union bound over the $N$ elements of $L$.
We now turn to the case where $x\in L$. If $v\in L\setminus\{x\}$, then $\deg_{G}(v)-1\sim\operatorname{Bin}\left(N,\left(1-\frac{D}{N}\right)\frac{1}{N}\right)$, which is stochastically bounded by $\operatorname{Bin}(N,1/N)$. Following the same steps as (8.4), we have that
$\mu_{x}^{\prime}\left(\deg_{G}(v)\geq\log^{2}n\right)\leq\mu_{x}^{\prime}\left(\deg_{G}(v)\geq\log^{2}N\right)\leq\exp\left(-\frac{2}{3}\log^{2}N\right).$
On the other hand, when $v=x$, $\deg_{G}(v)$ stochastically dominates $\operatorname{Bin}(N,D/N)$. By Bernstein’s inequality,
$\mu_{x}^{\prime}\left(\deg_{G}(v)\leq\frac{D}{2}\right)\leq\exp\left(-\frac{3}{56}D\right).$
It follows that
$\mu_{x}^{\prime}\left({\mathcal{H}}_{x}^{c}\right)$ $=\mu_{x}^{\prime}\left(\exists v\in L\setminus\{x\}:\deg_{G}(v)>\log^{2}n\text{ or }\deg_{G}(x)<\frac{D}{2}\right)$
$\leq\sum_{v\in L\setminus\{x\}}\mu_{x}^{\prime}\left(\deg_{G}(v)>\log^{2}n\right)+\mu_{x}^{\prime}\left(\deg_{G}(x)<\frac{D}{2}\right)$
$\leq N\exp\left(-\frac{2}{3}\log^{2}N\right)+\exp\left(-\frac{3}{56}D\right)=o(1).$ (8.6)
Together, (8.5), (8.6) and (8.3) prove the claim. ∎
8.1.2 Proof of Lemma 8.5
The advantage of using the auxiliary distributions is that they are product distributions conditioned on the infection times in $H$, which makes it easier to analyze the infection times in $[N]$. To see this concretely, let
$\mathbf{E}(j)=(E_{1}(j),\ldots,E_{K}(j))\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathrm{Exp}(1)^{\otimes K},\qquad j\in[N].$
Let us also define the collection of independent variables $\{U(j)\}_{j\in[N]}$ where $U(j)\sim\mathrm{Unif}(L)$ denotes the random neighbor of the vertex $j$ in the construction of $G\sim P_{\emptyset}^{\prime}$. We then have the representation
$\mathbf{T}(j)=\mathbf{T}(U(j))+\mathbf{E}(j),\qquad j\in[N].$ (8.7)
On the other hand, if $G\sim P_{v}^{\prime}$ for some $v\in L$, then we have the representation
$\mathbf{T}(j)=\begin{cases}\mathbf{T}(U(j))+\mathbf{E}(j)&\text{ with probability $1-D/N$}\\ \mathbf{T}(v)+\mathbf{E}(j)&\text{ with probability $D/N$.}\end{cases}$ (8.8)
We can also succinctly characterize the distributions of infection times in $[N]$ for the cases described in (8.7) and (8.8). Define the following (conditional) distributions over $\mathbf{t}:=(t_{1},\ldots,t_{K})\in\operatorname{\mathbb{R}}_{+}^{K}$,
$f_{v}(\mathbf{t})$ $:=\prod_{i=1}^{K}e^{-(t_{i}-T_{i}(v))}\mathbf{1}(t_{i}\geq T_{i}(v))=e^{-\sum_{i=1}^{K}t_{i}}\prod_{i=1}^{K}e^{T_{i}(v)}\mathbf{1}(t_{i}\geq T_{i}(v))$
$f_{\emptyset}(\mathbf{t})$ $:=\frac{1}{N}\sum_{w\in L}f_{w}(\mathbf{t})=\frac{e^{-\sum_{i=1}^{K}t_{i}}}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)}\mathbf{1}(t_{i}\geq T_{i}(w)).$
Then it is readily seen that
$\mathrm{Law}\left(\left.\{\mathbf{T}(j)\}_{j\in[N]}\right|\{\mathbf{T}(u)\}_{u\in L}\right)=\begin{cases}f_{\emptyset}^{\otimes N}&G\sim P_{\emptyset}^{\prime}\\ \left[\left(1-\frac{D}{N}\right)f_{\emptyset}+\frac{D}{N}f_{v}\right]^{\otimes N}&G\sim P_{v}^{\prime},v\in L.\end{cases}$ (8.9)
Our strategy moving forward is to determine when the distributions in (8.9) are indistinguishable, which amounts to bounding the total variation distance between them. To do so, we will instead bound a larger distance measure – the $\chi^{2}$ divergence. We recall that the $\chi^{2}$ divergence between probability measures $P$ and $Q$ is defined as
$\chi^{2}(P\|Q):=\int\frac{(dP-dQ)^{2}}{dQ}.$
The total variation distance and the $\chi^{2}$ divergence can be related through the inequality [polyanskiy_wu_information_theory, Proposition 7.15]
$\mathrm{TV}(P,Q)\leq 2\sqrt{\chi^{2}(P\|Q)}.$
For our setting, the $\chi^{2}$ divergence has two advantageous properties. First, it tensorizes; that is, for any positive integer $m$, it holds that [polyanskiy_wu_information_theory, Chapter 7.12]
$\chi^{2}\left(\left.P^{\otimes m}\right\|Q^{\otimes m}\right)\leq\left(1+\chi^{2}(P\|Q)\right)^{m}-1.$
Additionally, the formula for the $\chi^{2}$ divergence involves the difference between the densities of $P$ and $Q$, which makes it convenient to analyze mixture distributions, as is the case for our setting.
We apply these ideas to the conditional distributions in (8.9). Conditioned on $\{\mathbf{T}(u)\}_{u\in H}$, we have that
$\mathrm{TV}\left(f_{\emptyset}^{\otimes N},\left[\left(1-\frac{D}{N}\right)f_{\emptyset}+\frac{D}{N}f_{v}\right]^{\otimes N}\right)$ $\leq 2\sqrt{\chi^{2}\left(\left.\left[\left(1-\frac{D}{N}\right)f_{\emptyset}+\frac{D}{N}f_{v}\right]^{\otimes N}\right\|f_{\emptyset}^{\otimes N}\right)}$
$=2\sqrt{\left[1+\chi^{2}\left(\left.\left(1-\frac{D}{N}\right)f_{\emptyset}+\frac{D}{N}f_{v}\right\|f_{\emptyset}\right)\right]^{N}-1}$
$\leq 2\sqrt{\exp\left\{N\chi^{2}\left(\left.\left(1-\frac{D}{N}\right)f_{\emptyset}+\frac{D}{N}f_{v}\right\|f_{\emptyset}\right)\right\}-1}.$
In particular, we see that the conditional total variation tends to 0 if
$\chi^{2}\left(\left.\left(1-\frac{D}{N}\right)f_{\emptyset}+\frac{D}{N}f_{v}\right\|f_{\emptyset}\right)=o\left(\frac{1}{N}\right).$ (8.10)
We proceed by bounding the $\chi^{2}$ divergence between the two distributions. Letting $\mathbf{t}=(t_{1},\ldots,t_{K})\in\operatorname{\mathbb{R}}^{K}$, it holds that
[…]
To bound the $\chi^{2}$ divergence, we therefore need to study the likelihood ratio $f_{v}/f_{\emptyset}$. This is accomplished in the following result.
Lemma 8.7.
For every $v\in L$, it holds with probability $1-o(1)$ that
$\sup_{\mathbf{t}\in\operatorname{\mathbb{R}}_{+}^{K}}\frac{f_{v}(\mathbf{t})}{f_{\emptyset}(\mathbf{t})}\leq\log^{5K}n.$
We are now ready to complete the proof of the main result.
Proof of Lemma 8.5.
Let ${\mathcal{A}}_{v}$ denote the event described in Lemma 8.7; note that ${\mathcal{A}}_{v}$ is measurable with respect to $\{\mathbf{T}(u)\}_{u\in L}$. On this event, it follows from (8.11) that
[…]
In particular, for any $\epsilon>0$ the condition (8.10) is satisfied if
$K\leq\frac{1-2\alpha-\epsilon}{5}\left(\frac{\log n}{\log\log n}\right).$ (8.12)
We now translate our results on the conditional distributions $f_{\emptyset}^{\otimes N}$ and $[(1-D/N)f_{\emptyset}+(D/N)f_{v}]^{\otimes N}$ to their unconditional counterparts $P_{\emptyset}^{\prime}$ and $P_{v}^{\prime}$. If (8.12) holds, then
$\mathrm{TV}(P_{\emptyset}^{\prime},P_{v}^{\prime})$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\operatorname{\mathbb{E}}\left[\mathrm{TV}\left(f_{\emptyset}^{\otimes N},\left[\left(1-\frac{D}{N}\right)f_{\emptyset}+\frac{D}{N}f_{v}\right]^{\otimes N}\right)\right]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\operatorname{\mathbb{E}}\left[\mathrm{TV}\left(f_{\emptyset}^{\otimes N},\left[\left(1-\frac{D}{N}\right)f_{\emptyset}+\frac{D}{N}f_{v}\right]^{\otimes N}\right)\mathbf{1}({\mathcal{A}}_{v})\right]+\operatorname{\mathbb{P}}({\mathcal{A}}_{v}^{c})=o(1).$
Above, the inequality $(a)$ follows from the property that conditioning increases the total variation [polyanskiy_wu_information_theory, Theorem 7.5], and inequality $(b)$ holds since the total variation is always upper bounded by 1. Finally, $\mathrm{TV}(P_{\emptyset}^{\prime},P_{v}^{\prime})=o(1)$ implies that $\mathrm{TV}(P_{\emptyset},P_{v})=o(1)$ as well in light of (8.2). ∎
8.2 Analysis of the likelihood ratio: Proof of Lemma 8.7
In this section, we study the likelihood ratio
$\frac{f_{v}(\mathbf{t})}{f_{\emptyset}(\mathbf{t})}=\frac{\prod_{i=1}^{K}e^{T_{i}(v)}\mathbf{1}(t_{i}\geq T_{i}(v))}{\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)}\mathbf{1}(t_{i}\geq T_{i}(w))}.$
When there exists $i\in[K]$ such that $t_{i}<T_{i}(v)$, it is clear from the formula that the likelihood ratio is zero. On the other hand, when $t_{i}\geq T_{i}(v)$ for all $i\in[K]$,
$\frac{f_{v}(\mathbf{t})}{f_{\emptyset}(\mathbf{t})}$ $=\frac{\prod_{i=1}^{K}e^{T_{i}(v)}}{\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)}\mathbf{1}(t_{i}\geq T_{i}(w))}$
$\leq\frac{\prod_{i=1}^{K}e^{T_{i}(v)}}{\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))}$
$=\left(\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right)^{-1}.$
Above, the inequality on the second line holds since $\mathbf{1}(T_{i}(v)\geq T_{i}(w))\leq\mathbf{1}(t_{i}\geq T_{i}(w))$ if $t_{i}\geq T_{i}(v)$. Importantly, the upper bound derived for the likelihood ratio does not depend on $\mathbf{t}$. Hence, to control $\sup_{\mathbf{t}\in\operatorname{\mathbb{R}}_{+}^{K}}f_{v}(\mathbf{t})/f_{\emptyset}(\mathbf{t})$, we establish a lower bound for the quantity
$\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w)).$ (8.13)
We remark that the analysis of (8.13) is greatly facilitated by the way in which the graph $H$ is constructed (see Definition 8.2). To see this concretely, for each $v\in L$ let $v^{\prime}$ be the first vertex in $H^{\prime}$ on the unique path in $H$ connecting $v$ to $H^{\prime}$. Then we can write
$T_{i}(v)=T_{i}(v^{\prime})+X_{i}(v),\qquad i\in[K],$ (8.14)
where $X_{i}(v)$ is the time it takes for the cascade to spread from $v^{\prime}$ to $v$. Importantly, as the paths added to $H^{\prime}$ to form $H$ are all disjoint, $\{X_{i}(v)\}_{v\in L}$ is a collection of i.i.d. $\mathrm{Gamma}(\log^{5}N)$ random variables. Since the diameter of $H^{\prime}$ is assumed to be much smaller than $\log^{5}N$, we can show that $T_{i}(v^{\prime})=o(\log^{5}N)$, and the behavior of $T_{i}(v)$ is largely dictated by $X_{i}(v)$, a random variable that is independent across $v\in L$. As we shall see, this reduction to i.i.d. random variables leads to a tractable analysis of (8.13).
To formalize these ideas, we begin by defining the following “nice" event.
Definition 8.8.
Let $v\in L$. The event ${\mathcal{E}}_{v}$ holds if and only if the following conditions are satisfied:
  1. 1.
    $T_{i}(z)\leq\log^{2}N$ for all $z\in H^{\prime}$ and $i\in[K]$.
  2. 2.
    $T_{i}(v)\geq\frac{2}{3}\log^{5}N$ for all $i\in[K]$.
  3. 3.
    $\sum_{i=1}^{K}\left(X_{i}(v)-\log^{5}N\right)^{2}\leq 2K\log^{5}N$.
Importantly, as the following result shows, ${\mathcal{E}}_{v}$ holds with high probability.
Lemma 8.9.
For all $v\in L$, it holds that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v})=1-o(1)$.
Proof.
We let ${\mathcal{E}}_{v,1},{\mathcal{E}}_{v,2}$ and ${\mathcal{E}}_{v,3}$ denote the events corresponding to the first, second and third conditions of ${\mathcal{E}}_{v}$ in Definition 8.8. We start by showing that ${\mathcal{E}}_{v,1}$ holds with probability $1-o(1)$. By our assumption that the diameter of $H^{\prime}$ is at most $2\log N$ (see Definition 8.2), we have that $T_{i}(z)$ is stochastically upper bounded by a $\mathrm{Gamma}(2\log N,1)$ random variable. Hence
$\operatorname{\mathbb{P}}\left(T_{i}(z)\geq\log^{2}N\right)\leq e^{-\frac{1}{2}\log^{2}N}2^{\log N}\leq e^{-\frac{1}{4}\log^{2}N},$
where the first inequality is due to a Chernoff bound, and the second inequality holds for $N$ sufficiently large. Taking a union bound over all $z\in H^{\prime}$ and $i\in[K]$, we obtain
$\operatorname{\mathbb{P}}({\mathcal{E}}_{v,1}^{c})=\operatorname{\mathbb{P}}\left(\max_{z\in H^{\prime},i\in[K]}T_{i}(z)\geq\log^{2}N\right)\leq\sum_{z\in H^{\prime},i\in[K]}\operatorname{\mathbb{P}}\left(T_{i}(z)\geq\log^{2}N\right)\leq NKe^{-\frac{1}{4}\log^{2}N}=o(1).$ (8.15)
We now turn to the events ${\mathcal{E}}_{v,2}$ and ${\mathcal{E}}_{v,3}$. Notice that if ${\mathcal{E}}_{v,1}$ holds and there exists $i\in[K]$ such that $T_{i}(v)<\frac{2}{3}\log^{5}N$, then
$\sum_{i=1}^{K}\left(X_{i}(v)-\log^{5}N\right)^{2}\geq\left(\frac{\log^{5}N}{4}\right)^{2}>2K\log^{5}N,$ (8.16)
where the final inequality holds for $N$ sufficiently large. As a consequence of (8.16), we have that
${\mathcal{E}}_{v,1}\cap{\mathcal{E}}_{v,2}^{c}\subseteq{\mathcal{E}}_{v,3}^{c},$ (8.17)
so we focus on bounding the probability of the latter event. We do so through a second moment argument. As a shorthand, let $R:=\log^{5}N$. The first moment of each of the summands is $\operatorname{\mathbb{E}}[(X_{i}(v)-R)^{2}]=\mathrm{Var}(X_{i}(v))=R$, where we have used the fact that $X_{i}(v)\sim\mathrm{Gamma}(R,1)$. The second moment is
$\operatorname{\mathbb{E}}\left[(X_{i}(v)-R)^{4}\right]$ $=\operatorname{\mathbb{E}}[X_{i}(v)^{4}]-4R\operatorname{\mathbb{E}}[X_{i}(v)^{3}]+6R^{2}\operatorname{\mathbb{E}}[X_{i}(v)^{2}]-4R^{3}\operatorname{\mathbb{E}}[X_{i}(v)]+R^{4}$
$=R(R+1)(R+2)(R+3)-4R^{2}(R+1)(R+2)+6R^{3}(R+1)-4R^{4}+R^{4}$
$=3R^{2}+6R.$
In particular, $\mathrm{Var}((X_{i}(v)-R)^{2})=2R^{2}+6R\leq 8R^{2}$, and furthermore,
$\operatorname{Var}\left(\sum_{i=1}^{K}(X_{i}(v)-R)^{2}\right)=\sum_{i=1}^{K}\operatorname{Var}((X_{i}(v)-R)^{2})\leq 8KR^{2}.$
Finally, by Chebyshev’s inequality,
$\operatorname{\mathbb{P}}({\mathcal{E}}_{v,3}^{c})=\operatorname{\mathbb{P}}\left(\sum_{i=1}^{K}(X_{i}(v)-R)^{2}>2K\right)\leq\frac{1}{4K^{2}R^{2}}\operatorname{Var}\left(\sum_{i=1}^{K}(X_{i}(v)-R)^{2}\right)\leq\frac{2}{K}=o(1).$ (8.18)
Together, (8.15), (8.17) and (8.18) prove that $\operatorname{\mathbb{P}}({\mathcal{E}}_{v}^{c})=o(1)$ as claimed. ∎
Next, we state a simple consequence of Taylor’s theorem that will be useful for the proof of Lemma 8.7.
Lemma 8.10.
Suppose that $a,b$ are real numbers such that $b\neq 0$ and $a/b\geq-1/2$. Then
$e^{-a}\left(1+\frac{a}{b}\right)^{b}\geq e^{-2a^{2}/b}.$
Proof.
Taylor’s theorem implies that for any $z\geq-1/2$, it holds that $\log(1+z)\geq z-2z^{2}$. Using this inequality, we have that
$\log\left(e^{-a}\left(1+\frac{a}{b}\right)^{b}\right)$ $=-a+b\log\left(1+\frac{a}{b}\right)\geq-a+b\left(\frac{a}{b}-2\left(\frac{a}{b}\right)^{2}\right)=-\frac{2a^{2}}{b}.$
We now prove the main result of this section.
Proof of Lemma 8.7.
Throughout this proof, we shall condition on the infection times in $H^{\prime}$, given by the collection $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$, as well as $v$, given by $\mathbf{T}(v)$. We will also assume that the “nice" event ${\mathcal{E}}_{v}$ holds, which is measurable with respect to $\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)$. Finally, as a shorthand, we will denote $R:=\log^{5}N$.
Recall from Fact 8.3 that $\{\mathbf{T}(u)\}_{u\in L}$ is a collection of conditionally independent variables with respect to $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$. As a result, if we condition on both $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$ and $\mathbf{T}(v)$, then the terms
$\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w)),\qquad w\in L,$ (8.19)
are mutually (conditionally) independent. Moreover, each of the terms in (8.19) are bounded between 0 and 1. Hoeffding’s inequality therefore implies that
$\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\\ \geq\operatorname{\mathbb{E}}\left[\left.\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right|\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)\right]-N^{-1/3},$ (8.20)
with probability at least $1-2\exp(-2N^{1/3})$. We proceed by studying the conditional expectation in (8.20). To this end, for any $w\in L$, we denote $w^{\prime}$ to be the last vertex in $H^{\prime}$ on the unique path starting from $v_{0}$ and ending at $w$. We have that
$\operatorname{\mathbb{E}}\left[\left.\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right|\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)\right]$
$\hskip 56.9055pt=\frac{1}{N}\sum_{w\in L}\operatorname{\mathbb{E}}\left[\left.\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right|\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)\right]$
$\hskip 56.9055pt=\frac{1}{N}\sum_{w\in L}\operatorname{\mathbb{E}}\left[\left.\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right|\mathbf{T}(w^{\prime}),\mathbf{T}(v)\right]$
$\hskip 56.9055pt=\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}\operatorname{\mathbb{E}}\left[\left.e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right|T_{i}(w^{\prime}),T_{i}(v)\right].$ (8.21)
Above, the second equality uses that the conditional expectation depends only on the random variables $T_{i}(w),T_{i}(v)$, and that the distribution of $T_{i}(w)$ conditioned on $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$ is the same as the distribution conditioned on $T(w^{\prime})$ in light of (8.14). The final equality follows from the independence of the $T_{i}(w)$’s over $i\in[K]$.
We continue by lower bounding the conditional expectations in (8.21). Recall that for $w\in L$, we have from (8.14) that $T_{i}(w)=T_{i}(w^{\prime})+X_{i}(w)$ where $X_{i}(w)\sim\mathrm{Gamma}(R,1)$. Then
$\operatorname{\mathbb{E}}\left[\left.e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right|T_{i}(w^{\prime}),T_{i}(v)\right]$
$\hskip 56.9055pt=e^{T_{i}(w^{\prime})-T_{i}(v)}\operatorname{\mathbb{E}}\left[\left.e^{X_{i}(w)}\mathbf{1}(T_{i}(v)-T_{i}(w^{\prime})\geq X_{i}(w))\right|T_{i}(w^{\prime}),T_{i}(v)\right]$
[…]
$\hskip 56.9055pt=\frac{e^{T_{i}(w^{\prime})-T_{i}(v)}(T_{i}(v)-T_{i}(w^{\prime}))^{R}}{\Gamma(R+1)}$
$\hskip 56.9055pt\stackrel{{\scriptstyle(b)}}{{\geq}}\frac{e^{T_{i}(w^{\prime})-T_{i}(v)}(T_{i}(v)-T_{i}(w^{\prime}))^{R}}{3\sqrt{R}(R/e)^{R}}$
$\hskip 56.9055pt=\frac{1}{3\sqrt{R}}e^{-(T_{i}(v)-T_{i}(w^{\prime})-R)}\left(1+\frac{T_{i}(v)-T_{i}(w^{\prime})-R}{R}\right)^{R}$
$\hskip 56.9055pt\stackrel{{\scriptstyle(c)}}{{\geq}}\frac{1}{3\sqrt{R}}\exp\left\{-\frac{2(T_{i}(v)-T_{i}(w^{\prime})-R)^{2}}{R}\right\}$
$\hskip 56.9055pt\stackrel{{\scriptstyle(d)}}{{\geq}}\frac{1}{3\sqrt{R}}\exp\left\{-\frac{4}{R}(X_{i}(v)-R)^{2}-\frac{4}{R}\left(T_{i}(v^{\prime})+T_{i}(w^{\prime})\right)^{2}\right\}$
$\hskip 56.9055pt\stackrel{{\scriptstyle(e)}}{{\geq}}\frac{e^{-16}}{3\sqrt{R}}\exp\left\{-\frac{4}{R}(X_{i}(v)-R)^{2}\right\}.$ (8.22)
In the display above, the equality $(a)$ uses that $T_{i}(v)-T_{i}(w^{\prime})\geq 0$, which holds since $T_{i}(v)\geq 2R/3\geq\sqrt{R}\geq\max_{z\in G^{\prime}}T_{i}(z)$ on the event ${\mathcal{E}}_{v}$. The inequality $(b)$ is due to Stirling’s approximation. The inequality $(c)$ follows from Lemma 8.10, which can be applied here since $T_{i}(v)-T_{i}(w^{\prime})\geq R/2$ on ${\mathcal{E}}_{v}$. The inequality $(d)$ uses $(a+b)^{2}\leq 2a^{2}+2b^{2}$, and the inequality $(e)$ follows since $T_{i}(v^{\prime}),T_{i}(w^{\prime})\leq\sqrt{R}$ on the event ${\mathcal{E}}_{v}$.
Substituting the lower bound (8.22) into (8.21), we obtain
$\operatorname{\mathbb{E}}\left[\left.\frac{1}{|B|}\sum_{w\in B}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right|\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)\right]\\ \geq\left(\frac{e^{-16}}{3\sqrt{R}}\right)^{K}\exp\left\{-\frac{4}{R}\sum_{i=1}^{K}(X_{i}(v)-R)^{2}\right\}\stackrel{{\scriptstyle(f)}}{{\geq}}\left(\frac{e^{-30}}{\sqrt{R}}\right)^{K},$ (8.23)
where $(f)$ holds on the event ${\mathcal{E}}_{v}$.
Together, (8.20) and (8.23) imply that conditioned on $\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)$, and provided ${\mathcal{E}}_{v}$ holds, we have with probability at least $1-2\exp(2N^{1/3})$ that
$\frac{1}{|B|}\sum_{w\in B}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\geq\left(\frac{e^{-30}}{\sqrt{R}}\right)^{K}-N^{-1/3}\geq R^{-K}.$
Finally, the desired result holds since $\operatorname{\mathbb{P}}({\mathcal{E}}_{v})=1-o(1)$. ∎
Appendix A Proof of Proposition 4.1
It suffices to show that the model described in Proposition 4.1 satisfies the equation (4.1). As (4.1) describes a unique Markov process, the claim will then follow. We start by showing that it is unlikely for two vertices to become infected in a very small time interval.
Lemma A.1.
For any $t\geq 0$, it holds that
$\operatorname{\mathbb{P}}\left(\left.|I[t,t+\epsilon]|\geq 2\right|{\mathcal{F}}_{t}\right)=o(\epsilon).$
Proof.
Recall that for a subset of vertices $S\subset V$, we denote $\operatorname{\mathsf{cut}}(S):=\{(u,w)\in E(G):u\in S,v\notin S\}$. We also define $\tau_{1}$ and $\tau_{2}$ to be the first and second times after $t$ when a new vertex is added to the cascade, respectively. Conditionally on ${\mathcal{F}}_{t}$, the memoryless property of Exponential distributions implies that $\tau_{1}-t$ is equal in distribution to $\mathrm{Exp}(\lambda|\operatorname{\mathsf{cut}}(I[0,t])|)$, which stochastically dominates a $\mathrm{Exp}(\lambda n^{2})$ random variable (here, we have used that the number of edges in $G$ is at most $n^{2}$). By the memoryless property of Exponential distributions, $\tau_{2}-\tau_{1}$ is independent of $\tau_{1}-t$. Through similar argument as above, it holds that $\tau_{2}-\tau_{1}$ stochastically dominates a $\mathrm{Exp}(\lambda n^{2})$ random variable as well.
Putting everything together, the probability that at least two vertices join the cascade in the time interval $[t,t+\epsilon]$ is
$\operatorname{\mathbb{P}}(\tau_{2}-t\leq\epsilon|{\mathcal{F}}_{t})$ $\leq\operatorname{\mathbb{P}}\left(\tau_{1}-t\leq\epsilon\text{ and }\tau_{2}-\tau_{1}\leq\epsilon|{\mathcal{F}}_{t}\right)$
$=\operatorname{\mathbb{P}}\left(\tau_{1}-t\leq\epsilon|{\mathcal{F}}_{t}\right)\operatorname{\mathbb{P}}\left(\tau_{2}-\tau_{1}\leq\epsilon|{\mathcal{F}}_{t}\right)$
$\leq\left(1-e^{-\epsilon\lambda n^{2}}\right)^{2}$
$\leq\epsilon^{2}\lambda^{2}n^{4}=o(\epsilon).$
We now turn to the proof of the main result of this section.
Proof of Proposition 4.1.
We start by defining some relevant notation. Condition on ${\mathcal{F}}_{t}$, and let $v\in V\setminus I[0,t]$. Recalling that $\{F_{t}(e)\}_{e\in E(G)}$ is a collection of i.i.d. $\mathrm{Exp}(\lambda)$ random variables, we define
$W$ $:=\min\{F_{t}(w,v):w\in I[0,t],(w,v)\in E(G)\}.$
By basic properties of Exponential distributions, we have that $W\sim\mathrm{Exp}(\lambda|{\mathcal{N}}(v)\cap I[0,t]|)$. Additionally define the event ${\mathcal{A}}$ to hold if and only if at most one vertex joins the cascade in the interval $(t,t+\epsilon]$.
We claim that if ${\mathcal{A}}$ holds, then $v\in I[0,t+\epsilon]\iff W\leq\epsilon$. Indeed, if $W\leq\epsilon$, then from the description of $T(v)$ in the proposition statement, we have that $T(v)\leq t+\epsilon$, which implies that $v\in I[t,t+\epsilon]$. On the other hand, if $W>\epsilon$, then we consider two cases: (1) no vertex becomes infected in $[t,t+\epsilon]$ or (2) one vertex becomes infected in $[t,t+\epsilon]$. Note that no other case is possible if ${\mathcal{A}}$ holds. In the first case, it holds for all $e\in\operatorname{\mathsf{cut}}(I[0,t])$ that $F_{t}(e)>\epsilon$, hence the weight of any path starting in $I[0,t]$ and ending at $v$ is at least $\epsilon$, implying that $T(v)>t+\epsilon$. In the second case, if $v\in I[t,t+\epsilon]$ despite $W>\epsilon$, there must exist a path $P\in{\mathcal{P}}(v,t)$ of length at least 2 such that $\mathsf{weight}(P,t)\leq\epsilon$. However, this implies the existence of another vertex $u$ on the path which is not an element of $I[0,t]$. Since the portion of the path halting at $u$ has a smaller weight than $P$, it follows that $T(u)\in I[t,t+\epsilon]$. However, it is not possible for both $u$ and $v$ to be elements of $I[t,t+\epsilon]$ if ${\mathcal{A}}$ holds, hence we must have that $T(v)>t+\epsilon$.
As a result of the argument above, we have that
($\operatorname{\mathbb{P}}\left(\{v\in I[0,t+\epsilon]\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)$ =($=\operatorname{\mathbb{P}}\left(\{W\leq\epsilon\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)$
$=\operatorname{\mathbb{P}}(W\leq\epsilon|{\mathcal{F}}_{t})-\operatorname{\mathbb{P}}\left(\{W\leq\epsilon\}\cap{\mathcal{A}}^{c}|{\mathcal{F}}_{t}\right)$
$=\epsilon\lambda|{\mathcal{N}}(v)\cap I[0,t]|+o(\epsilon).$
Finally, since $\operatorname{\mathbb{P}}({\mathcal{A}}|{\mathcal{F}}_{t})=1-o(\epsilon)$, it holds that
$\operatorname{\mathbb{P}}\left(v\in I[0,t+\epsilon]|{\mathcal{F}}_{t}\right)=\epsilon\lambda|{\mathcal{N}}(v)\cap I[0,t]|+o(\epsilon),$
hence (4.1) is satisfied. ∎